### v2

# Coincidences of Division Fields of an elliptic curve defined over a number field

Abstract.
For an elliptic curve defined over a number field, the absolute Galois group acts on the group of torsion points of the elliptic curve, giving rise to a Galois representation in $\mathrm{GL}_{2}(\hat{\mathbb{Z}})$. The obstructions to the surjectivity of this representation are either local (i.e. at a prime), or due to nonsurjectivity on the product of local Galois images. In this article, we study an extreme case: the coincidence i.e. the equality of $n$-division fields, generated by the $n$-torsion points, attached to different positive integers $n$. We give necessary conditions for coincidences, dealing separately with vertical coincidences, at a given prime, and horizontal coincidences, across multiple primes, in particular when the Galois group on the $n$-torsion contains the special linear group. We also give a non-trivial construction for coincidences not occurring over $\mathbb{Q}$.
Key words and phrases:
Elliptic curves, Galois representations, Entanglement, Galois theory
1991 Mathematics Subject Classification:
Primary 11G05, 11F80; Secondary 11R32
Introduction
Let $F$ be a number field, $\overline{F}$ an algebraic closure of $F$ and $G_{F}=\operatorname{\mathrm{Gal}}(\overline{F}/F)$ its absolute Galois group. Let $E/F$ be an elliptic curve. We know that the absolute Galois group $G_{F}$ acts on the group $E_{\mathrm{tors}}$ of the torsion points of $E/F$ encoded by the Galois representation
$\rho_{E}:G_{F}\to\operatorname{\mathrm{Aut}}(E_{\mathrm{tors}})\simeq\mathrm{GL}_{2}(\hat{\mathbb{Z}}).$
Serre [Ser72] proved that, if $E/F$ does not have CM, then the image of $\rho_{E}$ has finite index in $\mathrm{GL}_{2}(\hat{\mathbb{Z}})$. If $E/F$ has CM, then there exists a choice of basis for $E_{\mathrm{tors}}$ such that $\rho_{E}(G_{F})$ is contained with finite index in a subgroup $N_{\delta,\phi}$, only depending on $\operatorname{\mathrm{End}}(E)$ as order in a quadratic field, see [LR22, Theorem 1.2]. Set $G=\mathrm{GL}_{2}(\hat{\mathbb{Z}})$ if $E/F$ does not have CM and $G=N_{\delta,\phi}$ if $E/F$ has CM. Let $G(p^{\infty})$, respectively $G(m)$, be the image of $G$ in $\mathrm{GL}_{2}(\mathbb{Z}_{p})$, respectively in $\mathrm{GL}_{2}(\mathbb{Z}/m\mathbb{Z})$. Therefore, the mod $m$ Galois representation
$\rho_{E,m}:G_{F}\to G(m)$
is surjective for all $m$ coprime to a fixed integer. For a positive integer $m$, the non surjectivity of $\rho_{E,m}$ can be explained by two phenomena:
  1. (1)
    Local conditions: The non surjectivity of the $p$-adic Galois representation $\rho_{E,p^{\infty}}:G_{F}\to G(p^{\infty})$ for some $p\mid m$.
  2. (2)
    Entanglement: The non surjectivity of $\rho_{E}(G_{F})$ on $\underset{\begin{subarray}{c}p\text{ prime}\\ p\mid m\end{subarray}}{\prod}\rho_{E,p^{\infty}}(G_{F})$.
If $F=\mathbb{Q}$, then $\rho_{E}$ is never surjective: we have at least that $\rho_{E,2^{\infty}}$ is not surjective or a Serre entanglement, see [DM22, Section 3.13.1].
Let $F(E[m])$, respectively $F(E[p^{\infty}])$, be the extension of $F$ generated by the coordinates of the $m$-torsion points of $E$, respectively the $p^{k}$-torsion points of $E$ for all $k$. The kernel of $\rho_{E,m}$ is $\operatorname{\mathrm{Gal}}(\overline{F}/F(E[m]))$. Entanglement is equivalent to the failure of the fields $\{F(E[p^{\infty}]):p\text{ prime},p\mid m\}$ to be linearly disjoint. In [CS23, Theorem 3.2] and [CP22, Theorem 1.1], Campagna, Pengo and Stevenhagen gave, for an elliptic curve $E/F$ with $\operatorname{\mathrm{End}}(E)\subset F$, an explicit finite set of primes $S$ such that
(*) $\operatorname{\mathrm{Gal}}(F(E_{\mathrm{tors}})/F)\simeq\operatorname{\mathrm{Gal}}(F(E[S^{\infty}])/F)\times\underset{\begin{subarray}{c}p\text{ prime}\\ p\notin S\end{subarray}}{\prod}\operatorname{\mathrm{Gal}}(F(E[p^{\infty}])/F)$
where $F(E[S^{\infty}])$ is the compositum of the $F(E[p^{\infty}])$ for $p\in S$. Let $\Delta_{F}$ be the discriminant of $F$, $\mathfrak{f}_{E}$ be the ideal conductor of $E$ and $\mathrm{N}(\mathfrak{f}_{E})$ be its norm. If $E/F$ does not have CM, the set $S$ consists of the primes $p$ satisfying at least one of the two following conditions:
  • $p\mid 2\cdot 3\cdot 5\cdot\Delta_{F}\cdot\mathrm{N}(\mathfrak{f}_{E})$,
  • $\rho_{E,p}$ is not surjective.
If $E/F$ has CM by $\mathcal{O}$, an order in $K$, the set $S$ consists of the primes dividing
$[\mathcal{O}_{K}:\mathcal{O}]\cdot\Delta_{F}\cdot\mathrm{N}(\mathfrak{f}_{E}).$
We describe what local conditions and entanglement imply about the division fields $F(E[m])$ for $m\geq 2$.
Local conditions. We know that if $p\geq 5$ and $\rho_{E,p}$ is surjective, then $\rho_{E,p^{\infty}}$ is surjective, see [KS09, Lemma 1]. But, if $\rho_{E,p}$ is not surjective, then we cannot deduce the image of $\rho_{E,p^{\infty}}$ from the image of $\rho_{E,p}$. It can be the full inverse image of $\rho_{E,p}(G_{F})$ in $\mathrm{GL}_{2}(\mathbb{Z}_{p})$, which is equivalent to having
$\operatorname{\mathrm{Gal}}(F(E[p^{k+1}]/F(E[p^{k}]))\simeq(\mathbb{Z}/p\mathbb{Z})^{4}$
for all $k\geq 1$, but it can be smaller.
Entanglement. We say that $E/F$ has an $(m,n)$-entanglement if
$F(E[m])\cap F(E[n])\neq F(E[\gcd(m,n)]).$
This corresponds to the linear dependance of the fields $F(E[m])$ and $F(E[n])$ over $F(E[\gcd(m,n)])$.
Our work focuses on the extreme case: a coincidence. We say that $E/F$ has an $(m,n)$-coincidence if $F(E[m])=F(E[n])$. Stevenhagen asked whenever we have a $(2^{k},2^{k+1})$-coincidence for an elliptic curve defined over $\mathbb{Q}$ and the answer was given by Rouse and Zureick-Brown in [RZB15, Remark 1.5]. Around the same time, in [BJ16], Brau and Jones gave a parametrization of elliptic curves $E/\mathbb{Q}$ such that $\mathbb{Q}(E[2])\subseteq\mathbb{Q}(E[3])$, which is equivalent to having a $(3,6)$-coincidence and corresponds to a $(2,3)$-entanglement. In [DLR23], Daniels and Lozano-Robledo ask when we have $\mathbb{Q}(\zeta_{p^{k}})\subseteq\mathbb{Q}(E[m])$ for a prime $p$ and, as a consequence of the Weil pairing, use it to study coincidence. In particular, they showed that the only possible $(m,n)$-coincidence with $m$ and $n$ prime powers are $(2,4)$ and $(2,3)$-coincidences.
In this article, we study the question of coincidence for elliptic curves over an arbitrary number field. As a first observation, we deduce from the isomorphism (*Introduction) the following result:
Lemma (Lemma 1.2).
Let $m$ and $n$ be two integers and set $m_{S}$, respectively $n_{S}$, to be the greatest divisor of $m$, respectively $n$, with only prime divisors in $S$. Suppose that $F(E[m])=F(E[n])$. Then
$F(E[m_{S}])=F(E[n_{S}])$
and
$\forall p\notin S,\quad F(E[p^{v_{p}(m)}])=F(E[p^{v_{p}(n)}]),$
where $v_{p}$ denotes the $p$-adic valuation.
We consider the explicit set $S$ given by Campagna et al. If $E/F$ does not have CM, then, for an $(m,n)$-coincidence, we have $v_{p}(m)=v_{p}(n)$ for all $p\notin S$. However, the set $S$ is not minimal for this property. In this article, we prove the following result:
Theorem (Theorem 3.14).
Let $E/F$ be an elliptic curve and $m,n\geq 1$. Suppose that $F(E[m])=F(E[n])$. Then, for all primes $p$ such that $v_{p}(m)\neq v_{p}(n)$, we have $p\mid 2\cdot\Delta_{F}\cdot\mathrm{N}(\mathfrak{f}_{E})$.
Corollary (Corollary 3.15).
Let $E/\mathbb{Q}$ be an elliptic curve, $\Delta_{E}$ be the minimal discriminant of $E$ and $m,n\geq 1$. Suppose that $\mathbb{Q}(E[m])=\mathbb{Q}(E[n])$. Then, for all primes $p$ such that $v_{p}(m)\neq v_{p}(n)$, we have $p\mid 2\Delta_{E}$.
If $m\mid n$, then $F(E[m])\subseteq F(E[n])$. For arbitrary positive integers $m$ and $n$, if a coincidence $F(E[m])=F(E[n])$ holds, then it remains true replacing $n$ by $\mathrm{lcm}(m,n)$. Thus, to obtain constraints on coincidences, it suffices to consider $m$ dividing $n$. Furthermore, we can reduce to the question of whether $F(E[m])=F(E[p^{k}m])$ for a prime $p$ and $k\geq 1$. Moreover, considering a set $S$ satisfying the isomorphism (*Introduction), it suffices to consider $m$ with only prime divisors in $S\cup\{p\}$. Then, we consider the following guiding question:
Question.
Let $p$ be a prime, $k\geq 1$ and $m$ be an integer with only prime divisors in $S\cup\{p\}$. When do we have $F(E[m])=F(E[p^{k}m])$?
We can reformulate this question, considering $p\nmid m$ and the following situations:
  1. -
    Horizontal coincidences
    • $F(E[m])=F(E[pm])$ for some $k\geq 1$
  2. -
    Vertical coincidences
    • $F(E[m])\neq F(E[pm])=\dots=F(E[p^{k}m])$ for some $k\geq 2$, or
    • $F(E[2m])\neq F(E[4m])=\dots=F(E[2^{k}m])$ for some $k\geq 3$.
We know by Theorem 3.20 that there are no other cases. In addition to refining the set of possible prime divisors as seen above, we give constraints on the possible exponents on the prime divisors. The two following theorems concern horizontal coincidences.
Theorem (Corollary 1.6).
Let $m,n\geq 3$, $p$ be a prime such that $p>q$ for all primes $q\mid m$. Suppose that $\mathbb{Q}(\zeta_{p^{v_{p}(n)}})\not\subseteq F$. If $E/F$ has an $(m,n)$-coincidence, then $v_{p}(n)=1$.
Theorem (Corollary 2.9 and Remark 2.10).
Let $E/F$ be an elliptic curve with an $(m,n)$-coincidence. Let $p$ be a prime such that $p\mid n$ and $p\nmid m$. Let $\mathfrak{p}$ an ideal above $p$ such that the index of ramification of $F/\mathbb{Q}$ at $\mathfrak{p}$ is prime to $\varphi(p^{k})$. Then we are in one of the following cases:
  • $v_{p}(n)=1$ and $E/F$ has bad reduction at $\mathfrak{p}$,
  • $v_{p}(n)=2$, $p=2$ and at $\mathfrak{p}$, $E/F$ has either additive or non split multiplicative reduction,
  • $v_{p}(n)=2$, $p=3$, and $E/F$ has additive and potential good reduction at $\mathfrak{p}$,
  • $v_{p}(n)=3$ or $4$, $p=2$ and $E/F$ has additive and potential good reduction at $\mathfrak{p}$.
We next consider vertical coincidences. The following theorem holds without any hypothesis either on the curve $E/F$ or the field $F$.
Theorem (Theorem 3.20 and Proposition 3.16).
Let $q=p$ and $k\geq 1$ if $p$ is odd, or $q=p^{2}$ and $k\geq 2$ if $p$ is even. Let $E/F$ be an elliptic curve. If $F(E[p^{k}])=F(E[p^{k+1}])$, then $F(E[q])=F(E[p^{k+1}])$ and $p^{4}\mid[G:\rho_{E}(G_{F})]$.
If $F=\mathbb{Q}$, we know by [DLR23, Theorem 1.4] that a $(p^{k},p^{k+1})$-coincidence is possible only for $p=2$. We prove that, more generally, this is true if $F\cap\mathbb{Q}(\zeta_{p^{k}})=\mathbb{Q}$, this is Corollary 3.10. If $E$ has CM by a quadratic field $K$ and $F\subseteq K(j(E))$, then Proposition 3.38 is more precise: a $(p^{k},p^{k+1})$-coincidence is not possible for $p$ odd and $k\geq 2$.
Furthermore, we observe that, if $m\mid n$, then the reduction map $\rho_{E,n}(G_{F})\to\rho_{E,m}(G_{F})$ is an isomorphism. This motivates the study of coincidences in chains
$F(E[m])=F(E[pm])=\dots=F(E[p^{k}m])$
as formulated in the guiding question, and the associated problem of splittings of the surjections $\rho_{E,pm}(G_{F})\to\rho_{E,m}(G_{F})$ introduced in Subsection 3.4. We prove the following result:
Theorem (Theorem 3.35).
Let $p$ be a prime. Let $q=p$ if $p\neq 2,3$, or $q=p^{2}$ if $p=2,3$. If $\rho_{E,q}(G_{F})$ contains a conjugates of $\begin{pmatrix}1&1\\ 0&1\end{pmatrix}$, then $F(E[q])\neq F(E[pq])$.
The next result is both valid for both horizontal and vertical coincidences. Obviously, an $(m,n)$-coincidence is not possible if $\rho_{E,m}$ and $\rho_{E,n}$ are both surjective, since $\mathrm{GL}_{2}(m)$ and $\mathrm{GL}_{2}(n)$ are not isomorphic for $m\neq n$. Daniels and Lozano-Robledo compared the abelian part of the division field to show that $E/\mathbb{Q}$ does not have $(m,n)$-coincidence if only $\rho_{E,m}$ is surjective. We use the same idea to show a similar result in case where $\rho_{E,m}$ is large.
Theorem (Theorem 4.4).
Let $m$ be an odd integer and $n\nmid m$ such that $\zeta_{n}\notin F$. Suppose that $\rho_{E,m}(G_{F})$ contains $\mathrm{SL}_{2}(\mathbb{Z}/m\mathbb{Z})$ and that $E/F$ has an $(m,n)$-coincidence. Then
$3\mid m,\quad\text{and}\quad\mathrm{D}(\rho_{E,m}(G_{F}))\neq\mathrm{SL}_{2}(\mathbb{Z}/m\mathbb{Z}),\quad\text{and}\quad F(\zeta_{n})\subseteq L$
with $L$ a $\mathbb{Z}/3\mathbb{Z}$-extension of $F(\zeta_{m})$.
Up to this point, we have only examined necessary condition for having a coincidence. We end the presentation of the result in this article with the statement of sufficient conditions. Over $\mathbb{Q}$, the only possible $(p^{k},p^{k+1})$-coincidence for a prime $p$ is a $(2,4)$-coincidence. It is even conjectured that the known $(2,4)$, $(2,3)$, $(2,6)$ and $(3,6)$-coincidences are the only possible coincidences over $\mathbb{Q}$. For any elliptic curve $E/F$, we can construct an $(m,n)$-coincidence with a base change from $F$ to $F(E[\mathrm{lcm}(m,n)])$, but such a base change provides a trivial construction. In Section 3 we prove the existence of a $(4,8)$-coincidence with a base change from $\mathbb{Q}$ to an extension linearly disjoint from $\mathbb{Q}(E[4])$:
Theorem (Theorem 3.1).
There are infinitely many isomorphism classes of elliptic curves $E/\mathbb{Q}$ such that there exists a number field $L$ with Galois group $(\mathbb{Z}/2\mathbb{Z})^{r}$ over $\mathbb{Q}$ with $1\leq r\leq 4$ satisfying
$L(E[4])=L(E[8])\neq L.$
Together with this theorem, in Subsection 3.1 we define the notion of a minimal base change and give necessary and sufficient condition to construct an $(m,mn)$-coincidence by a minimal base change of the ground field.
The structure of the paper is as follows. Section 1 contains a useful consequence of the Weil pairing and preliminary constraints for a coincidence. Section 2 deals with horizontal coincidences using the link between the ramification of $F(E[m])/F$ and the reduction type of $E/F$. Section 3 focuses on vertical coincidences. Section 4 concerns the case where $\rho_{E,m}(G_{F})$ contains $\mathrm{SL}_{2}(\mathbb{Z}/m\mathbb{Z})$. In Section 5, we use results of previous sections to give obstructions to a coincidence $F(E[m])=F(E^{\prime}[n])$ where $E/F$ and $E^{\prime}/F$ are different elliptic curves.
Acknowledgement
I thank my supervisors Samuele Anni and David Kohel for their many useful and interesting suggestions. I also thank Luis Dieulefait for discussion about constructing a non-trivial coincidence over a number field which does not happen over $\mathbb{Q}$.
1. Notations and preliminaries
Through this article we will use the following notations:
    For a set $S$, we denote by #$\#S$ its cardinality.
    For a positive integer $m$ and a prime $p$, $v_{p}(m)$ denotes the valuation of $m$ at $p$.
    For a positive integer $m$, we set $\mathrm{GL}_{2}(m):=\mathrm{GL}_{2}(\mathbb{Z}/m\mathbb{Z})$ and $\mathrm{SL}_{2}(m):=\mathrm{SL}_{2}(\mathbb{Z}/m\mathbb{Z})$.
    For finite extensions $K\subseteq L\subseteq M$ such that $M/L$ is Galois, and a prime ideal $\mathfrak{p}$ of $L$, we denote by $e_{\mathfrak{p}}(M/K)$ the index of ramification of $M/K$ at $\mathfrak{p}$.
    $F$ is a number field, $\overline{F}$ denotes an algebraic closure of $F$ and $G_{F}=\operatorname{\mathrm{Gal}}(\overline{F}/F)$ its absolute Galois group.
    $(\zeta_{m})_{m\geq 1}$ is a compatible system of primitive $m$-th root of unity in $\overline{F}$, that is $\zeta_{mn}^{n}=\zeta_{m}$ for $m,n\geq 1$.
    $\Delta_{F}$ denotes the discriminant of the number field $F$.
    $\mathcal{O}_{F}$ denotes the ring of integers of $F$.
First of all, we prove Proposition 1.2 given in the introduction after some preliminaries. Let $E/F$ be an elliptic curve. If $E/F$ has CM, we suppose that $F$ contains the CM field. From Serre’s open image theorem [Ser72, Section 4.4, Theorem 3’], there exists a finite set $S$ of rational primes such that
(1) $\operatorname{\mathrm{Gal}}(F(E_{\mathrm{tors}})/F)\simeq\operatorname{\mathrm{Gal}}(F(E[S^{\infty}])/F)\times\underset{\begin{subarray}{c}p\text{ prime}\\ p\notin S\end{subarray}}{\prod}\operatorname{\mathrm{Gal}}(F(E[p^{\infty}])/F)$
where $F(E[S^{\infty}])$ is the compositum of $F(E[p^{\infty}])$ for $p\in S$. The set $S$ is not unique and we are interested in finding the smallest possible set $S$. As underlined in the introduction, Campagna, Pengo and Stevenhagen gave a possible choice of $S$, distinguishing CM and non-CM case. All tensor products are taken over the base field $F$.
Remark 1.1.
Let $I$ be a countable set, $(L_{i})_{i\in I}$ be a family of linear disjoint extensions of $F$ and $K/F$ be a subfield of $\underset{i\in I}{\bigotimes}L_{i}$. Suppose that $K=\underset{i\in I}{\bigotimes}K_{i}$ with $K_{i}\subseteq L_{i}$ for all $i\in I$. For all $i\in I$, we have $K_{i}\subseteq K\cap L_{i}$ and
$\underset{i\in I}{\bigotimes}K_{i}\subseteq\underset{i\in I}{\bigotimes}(K\cap L_{i})\subseteq K=\underset{i\in I}{\bigotimes}K_{i}.$
Hence $K_{i}=K\cap L_{i}$. We recall that $\bigotimes K_{i}$ is a field if and only if the extensions $K_{i}/F$ are linearly disjoint, which means by definition that the surjection $\bigotimes K_{i}\to\prod K_{i}$ on the compositum is an isomorphism.
We deduce the following proposition, which does not depends on the set $S$. For an integer $m$, let $\gcd(m,S)$ be the greatest divisor of $m$ with only prime divisors in $S$.
Lemma 1.2.
Let $m$ and $n$ be two positive integers and suppose that $E/F$ has an $(m,n)$-coincidence. Then
$F(E[\gcd(m,S)])=F(E[\gcd(n,S)])$
and
$\forall p\notin S,\quad F(E[p^{v_{p}(m)}])=F(E[p^{v_{p}(n)}]).$
Proof.
From the isomorphism (1), we know that $F(E_{\mathrm{tors}})=\underset{i\in I}{\bigotimes}F(E[i^{\infty}])$ for
$I=\{S\}\cup\{p:p\text{ is prime and }p\notin S\}.$
By linear independance of Galois extensions, we have the following decompositions
$F(E[m])=F(E[\gcd(m,S)])\cdot\underset{p\notin S}{\prod}F(E[p^{v_{p}(m)}]),$
$F(E[n])=F(E[\gcd(n,S)])\cdot\underset{p\notin S}{\prod}F(E[p^{v_{p}(n)}]).$
By Remark 1.1 these decompositions are unique, therefore the proposition holds true. ∎
We will rely on this consequence of the Weil pairing:
Proposition 1.3.
Let $E/F$ be an elliptic curve and $m$ be a positive integer. The division field $F(E[m])$ contains $F(\zeta_{m})$. In particular, the image of $\det\circ\rho_{E,m}$ is equal to the image of $\operatorname{\mathrm{Gal}}(F(\zeta_{m})/F)$ in $(\mathbb{Z}/m\mathbb{Z})^{*}$.
Proof.
The inclusion $F(\zeta_{m})\subseteq F(E[m])$ follows from the Galois invariance of the Weil pairing, see [Sil09, Corollary 8.1.1]. Let $e_{m}$ be the Weil pairing on $E[m]$ and $(P,Q)$ be a basis of the $m$-torsion such that $e_{m}(P,Q)=\zeta_{m}$. Now, for $\sigma\in\operatorname{\mathrm{Gal}}(F(E[m])/F)$, the Galois invariance of $e_{m}$ gives
$e_{m}(\rho_{E,m}(\sigma)(P,Q))=e_{m}(\sigma(P),\sigma(Q))=\sigma(e_{m}(P,Q))=\sigma(\zeta_{m}),$
and the Weil pairing being bilinear and alternating, we have
$e_{m}(\rho_{E,m}(\sigma)(P,Q))=e_{m}(P,Q)^{\det\circ\rho_{E,m}(\sigma)}=\zeta_{m}^{\det\circ\rho_{E,m}(\sigma)}.$
The result follows. ∎
Proposition 1.3 implies that, if $F(E[n])=F(E[m])$ for some $n\geq m$, then $F(\zeta_{n})\subseteq F(E[m])$. A recurring strategy will be to give restrictions on having this inclusion.
Remark 1.4.
Let $n$ and $m$ be two integers such that $m<n$. Then there exists a prime $p$ such that $v_{p}(n)-v_{p}(m)=k\geq 1$. On the one hand, we have
$F(E[m])\subseteq F(E[p^{k}m])\subseteq F(E[\mathrm{lcm}(m,n)])=F(E[m])F(E[n]).$
On the other hand, we have $F(\zeta_{p^{k}m})\subseteq F(E[p^{k}m])$. Therefore, each time we have $F(\zeta_{p^{k}m})\not\subseteq F(E[m])$ for some $k\geq 1$, it follows that $F(E[m])\neq F(E[n])$ for all $n$ such that $v_{p}(n)-v_{p}(m)\geq k$.
As a first attempt, we investigate the possibility of an $(m,n)$-coincidence, simply by using the inclusions of fields $F(\zeta_{m})\subseteq F(E[m])$ and of groups $\rho_{E,m}(G_{F})\leq\mathrm{GL}_{2}(m)$, and the resulting divisibility of degrees and orders. The next proposition tells us that, if $F(E[n])=F(E[m])$, subject to an additional condition on $F$, then the primes greater than every prime dividing $m$ can divide $n$ to at most power $1$, unless $m$ is a power of $2$, in which case $3$ can divide $n$ with possibly a greater power than $1$.
Proposition 1.5.
Let $m\geq 2$, $p$ be a prime such that $p>q$ for all primes $q\mid m$ and $r$ be the largest integer such that $\mathbb{Q}(\zeta_{p^{r}})\subseteq F\cap\mathbb{Q}(\zeta_{p^{\infty}})$. Let $E/F$ be an elliptic curve such that $F(\zeta_{p^{k}})\subseteq F(E[m])$ with $k>r$. Then, $k=1$ (and $r=0$), unless $(m,p)=(2^{j},3)$ for some $j\geq 1$, in which case either $r=0$ and $k\leq 2$, or $r=k-1$.
Proof.
Suppose that $r>0$, or $r=0$ and $k\geq 2$. We will prove that $(m,p)=(2^{j},3)$, and $r=k-1$ or $r=0$ and $k=2$. By assumption, $p^{k-r}\mid[F(\zeta_{p^{k}}):F]$ if $r>0$ or $p^{k-1}\mid[F(\zeta_{p^{k}}):F]$ if $r=0$. In any case, $p$ divides $[F(\zeta_{p^{k}}):F]$. Since $F(\zeta_{p^{k}})\subseteq F(E[m])$, we have
$[F(\zeta_{p^{k}}):F]\bigm{|}[F(E[m]):F]\bigm{|}\#\mathrm{GL}_{2}(m),$
and
#$\#\mathrm{GL}_{2}(m)=\underset{\begin{subarray}{c}q^{j}\mid m\\ j=v_{q}(m)\end{subarray}}{\prod}\#\mathrm{GL}_{2}(q^{j})=\underset{\begin{subarray}{c}q^{j}\mid m\\ j=v_{q}(m)\end{subarray}}{\prod}q^{4(j-1)+1}(q-1)^{2}(q+1).$
Therefore, since $q<p$ for all $q\mid m$, we obtain $p=q+1$ for some $q$ dividing $m$ and so $m=2^{j}$ for some $j\geq 1$ and $p=3$. In this case, $k-r=1$ if $r>0$ and $k-1=1$ if $r=0$. ∎
Corollary 1.6.
Under the hypotheses of Proposition 1.5, let $n$ be an integer such that $v_{p}(n)=k$ and suppose that $E/F$ has an $(m,n)$-coincidence. Then, $k=1$, unless $(m,p)=(2^{j},3)$ for some $j\geq 1$, in which case $r=0$ and $k\leq 2$, or $r=k-1$.
In Corollary 2.9, in the next section, we extend Proposition 1.5 by replacing "$q<p$ for all $q\mid m$" by "$p\nmid m$", at the expense of adding conditions on the ramification at $p$ in $F$ or on the reduction type of $E$ at $p$.
2. Horizontal coincidence : ramification behaviour
We talk about horizontal coincidence if we have an $(m,n)$-coincidence and the sets of prime divisors of $m$ and $n$ are not the same. In this section, we study the obstructions to horizontal coincidences given by the type of reduction of the elliptic curve and the resulting ramification.
2.1. Ramification and reduction type
Let $\mathfrak{p}$ be a prime of $\mathcal{O}_{F}$, whose residue characteristic is $p$. We recall the criterion of Néron-Ogg-Shafarevitch:
Proposition 2.1 ([Sil09, VII, Theorem 7.1]).
Let $E/F$ be an elliptic curve. If $E/F$ has good reduction at $\mathfrak{p}$, then $F(E[m])/F$ is unramified at $\mathfrak{p}$ for all $m$ such that $p\nmid m$.
Moreover, the theory of Tate curves gives constraints on the ramification when the reduction is multiplicative:
Proposition 2.2.
Let $E/F$ be an elliptic curve and $m\geq 2$ such that $p\nmid m$. If $E/F$ has split multiplicative reduction at $\mathfrak{p}$ or if $E/F$ has multiplicative reduction $\mathfrak{p}$ and $p$ is odd, then $F(E[m])/F$ is tamely ramified at $\mathfrak{p}$. If $E/F$ has non split multiplication at $\mathfrak{p}$ and $p$ is even, then $v_{p}(e_{\mathfrak{p}}(F(E[m])/F))\leq 1$.
Proof.
First, we suppose that $E/F$ has split multiplicative reduction at $\mathfrak{p}$. Let $F_{\mathfrak{p}}$ be the completion of $F$ at $\mathfrak{p}$. We have, from [Sil94, V.Theorem 5.3], that $E$ is isomorphic over $F_{\mathfrak{p}}$ to the Tate curve $E_{q}$ for some $q\in F_{\mathfrak{p}}^{*}$ (for the definition of $E_{q}$, see [Sil94, V.Theorem 3.1]). We consider the $\mathfrak{p}$-adic uniformization:
$\overline{F_{\mathfrak{p}}}^{*}/q^{\mathbb{Z}}\overset{\sim}{\longrightarrow}E_{q}(\overline{F_{\mathfrak{p}}}).$
Restricting to the group of $m$-torsion on each side, we obtain an isomorphism
$\phi:\left(\zeta_{m}^{\mathbb{Z}}Q^{\mathbb{Z}}\right)/q^{\mathbb{Z}}\overset{\sim}{\longrightarrow}E_{q}[m],$
where $Q=q^{\frac{1}{m}}$ is a $m$-th root of $q$. The action of $\operatorname{\mathrm{Gal}}(\overline{F_{\mathfrak{p}}}/F_{\mathfrak{p}})$ on $E_{q}[m]$ is compatible with its action on $\left(\zeta_{m}^{\mathbb{Z}}Q^{\mathbb{Z}}\right)/q^{\mathbb{Z}}$ (see [Sil09, V, Theorem 5.3]). Let $I_{\mathfrak{p}}$ be the inertia subgroup of $\operatorname{\mathrm{Gal}}(\overline{F_{\mathfrak{p}}}/F_{\mathfrak{p}})$ and let $\sigma\in I_{\mathfrak{p}}$. Since $p\nmid m$, the extension $F_{\mathfrak{p}}(\zeta_{m})/F_{\mathfrak{p}}$ is unramified, and so $\sigma(\zeta_{m})=\zeta_{m}$. Since $Q$ is a root of $X^{m}-q$, so is $\sigma(Q)$. Therefore, there exists $a\in\mathbb{Z}/m\mathbb{Z}$ such that $\sigma(Q)=\zeta_{m}^{a}Q$. We set $P_{1}=\phi(\zeta_{m})$ and $P_{2}=\phi(Q)$. Then
$\sigma(P_{1})=\sigma(\phi(\zeta_{m}))=\phi(\sigma(\zeta_{m}))=\phi(\zeta_{m})=P_{1}$
and,
$\sigma(P_{2})=\sigma(\phi(Q))=\phi(\sigma(Q))=\phi(\zeta_{m}^{a}Q)=a\phi(\zeta_{m})+\phi(Q)=aP_{1}+P_{2}.$
Hence, for all $\sigma\in I_{\mathfrak{p}}$, there exists $a\in\mathbb{Z}/m\mathbb{Z}$ such that
$\rho_{E,p}(\sigma)=\begin{pmatrix}1&a\\ 0&1\end{pmatrix}.$
It follows that the image of the wild inertia by $\rho_{E,p}$ is included in a group of order $m$. However, as observed by Serre in [Ser72, Section 1.1], $I_{\mathfrak{p}}$ is a pro-$p$-group, and so its image by $\rho_{E,p}$ is a $p$-group. So it is trivial. Hence, $F(E[m])/F$ is tamely ramified at $\mathfrak{p}$.
Now, suppose that $E/F$ has non split multiplicative reduction at $\mathfrak{p}$, and let $L/F$ be the quadratic extension where the reduction is split. Then $L(E[m])/L$ is tamely ramified at $p$. Moreover,
$e_{\mathfrak{p}}(L(E[m])/F)=e_{\mathfrak{p}}(L(E[m])/L)e_{\mathfrak{p}}(L/F)$
and so
$e_{\mathfrak{p}}(F(E[m])/F)\mid e_{\mathfrak{p}}(L(E[m])/F)\mid 2e_{\mathfrak{p}}(L(E[m])/L),$
which completes the proof. ∎
Finally, we also have constraints on the ramification in case of additive reduction:
Proposition 2.3.
Let $E/F$ be an elliptic curve and $m\geq 2$ such that $p\nmid m$. If $p>3$ and $E/F$ has additive reduction at $\mathfrak{p}$, or if $p=3$ and $E/F$ does not have potential good reduction at $\mathfrak{p}$, then $F(E[m])/F$ is tamely ramified at $\mathfrak{p}$.
Proof.
If $E/F$ has potential good reduction, the proposition follows from [ST68, Section 2, Corollary 2]. If $E/F$ does not have potential good reduction, then the results follows from Proposition 2.2 and [Sil09, Appendix C, Theorem 14.1], since a quadratic extension cannot be widely ramified outside $2$. ∎
Finally, let us recall the following result:
Proposition 2.4 ([Ann14, Section 4.2]).
Let $E/F$ be an elliptic curve and $m\geq 2$ such that $p\nmid m$. Suppose that $E/F$ has additive reduction at $\mathfrak{p}$. There exists an extension $L/F$ of degree dividing $24$ such that $E/L$ has stable reduction at $\mathfrak{p}$.
2.2. Ramification and entanglement
Let $p$ be a prime and $\mathfrak{p}$ be a prime ideal of $F$ above $p$. Set $e=e_{\mathfrak{p}}(F/\mathbb{Q})$ the ramification index of $\mathfrak{p}$ in $F/\mathbb{Q}$. We know that, if $F(E[n])\subseteq F(E[m])$ then $F(\zeta_{p^{k}})\subseteq F(E[m])$ for all $p^{k}\mid n$. In particular, $e_{\mathfrak{p}}(F(\zeta_{p^{k}})/F)$ divides $e_{\mathfrak{p}}(F(E[m])/F)$ for all $p^{k}\mid n$. Lemma 2.5 gives information about $e_{\mathfrak{p}}(F(\zeta_{p^{k}})/F)$.
The map $\varphi:\mathbb{Z}\to\mathbb{Z}$ denotes the Euler totient function.
Lemma 2.5.
We have $v_{p}(e)\geq k-1-v_{p}(e_{\mathfrak{p}}(F(\zeta_{p^{k}})/F))$. Moreover, if $e_{\mathfrak{p}}(F(\zeta_{p^{k}})/F)=1$, then $\varphi(p^{k})\mid e$.
Proof.
The extension $F(\zeta_{p^{k}})/F$ is Galois and so the index of ramification above $\mathfrak{p}$ only depends on $\mathfrak{p}$. We have
$e_{\mathfrak{p}}(F(\zeta_{p^{k}})/F)e$ $=e_{\mathfrak{p}}(F(\zeta_{p^{k}})/\mathbb{Q})$
$=e_{\mathfrak{p}}(F(\zeta_{p^{k}})/\mathbb{Q}(\zeta_{p^{k}}))e_{\mathfrak{p}}(\mathbb{Q}(\zeta_{p^{k}})/\mathbb{Q})$
$=e_{\mathfrak{p}}(F(\zeta_{p^{k}})/\mathbb{Q}(\zeta_{p^{k}}))\varphi(p^{k})\cdot$
Since $v_{p}(\varphi(p^{k}))=k-1$ we obtain the first statement. The second follows from the previous equality. ∎
The previous section gives information about $e_{\mathfrak{p}}(F(E[m])/F)$, summarized in Theorem 2.6. These results give constraints on having an $(m,n)$-coincidence when $m$ and $n$ do not have the same prime divisors.
Theorem 2.6.
Let $m\geq 2$ such that $p\nmid m$. Let $E/F$ be an elliptic curve. The valuation at $\mathfrak{p}$ of the ramification index $e_{\mathfrak{p}}(F(E[m])/F)$ appears in the table below together with sufficient conditions on the reduction of $E/F$ at $\mathfrak{p}$.
Sufficient condition on $E/F$ $t=e_{\mathfrak{p}}(F(E[m])/F)$
good reduction at $\mathfrak{p}$ $t=1$
multiplicative red. at $\mathfrak{p}$ with $p$ odd
additive red. at $\mathfrak{p}$, $p>3$ $v_{p}(t)=0$
additive, not potentially good red. at $\mathfrak{p}$ with $p=3$
multiplicative red. at $\mathfrak{p}$ with $p=2$ $v_{p}(t)\leq 1$
additive, potentially good red. at $\mathfrak{p}$ with $p=3$
additive red. at $\mathfrak{p}$ with $p=2$ $v_{p}(t)\leq 3$
Proof.
Here is the table, with an additional column with the propositions required for the proof.
Sufficient condition on $E/F$ $t=e_{\mathfrak{p}}(F(E[m])/F)$ Proof
good red. at $\mathfrak{p}$ $t=1$ Proposition 2.1
mult. red. at $\mathfrak{p}$ with $p$ odd Proposition 2.2
add. red. at $\mathfrak{p}$, $p>3$ $v_{p}(t)=0$ Proposition 2.3
add., no pot. good red. at $\mathfrak{p}$ with $p=3$ Proposition 2.3
mult. red. at $\mathfrak{p}$ with $p=2$ $v_{p}(t)\leq 1$ Proposition 2.2
add., pot. good red. at $\mathfrak{p}$ with $p=3$ Proposition 2.4
add. red. at $\mathfrak{p}$ with $p=2$ $v_{p}(t)\leq 3$ Proposition 2.4
Remark 2.7.
By Lemma 2.5, we obtain the table below, in which we present the necessary condition on the ramification of $F/\mathbb{Q}$ to obtain the index of ramification as in previous theorem.
$s=e_{\mathfrak{p}}(F(\zeta_{p^{k}})/F)$ Necessary condition on $F/\mathbb{Q}$
$s=1$ $\varphi(p^{k})\mid e$
$v_{p}(s)=0$ $v_{p}(e)\geq k-1$
$v_{p}(s)\leq 1$ $v_{p}(e)\geq k-2$
$v_{p}(s)\leq 3$ $v_{p}(e)\geq k-4$
With the notation of Theorem 2.6 and Remark 2.7, if $F(\zeta_{p^{k}})\subseteq F(E[m])$, then we must have $s\mid t$. Therefore, the tables give restrictions on having $F(\zeta_{p^{k}})\subseteq F(E[m])$. For example, if we have this inclusion and $E/F$ has good reduction at $\mathfrak{p}$, then $\varphi(p^{k})$ must divide the index of ramification of $F/\mathbb{Q}$ at $\mathfrak{p}$. In the following corollary, we consider the case of $F/\mathbb{Q}$ unramified above $p$.
Corollary 2.8.
Let $E/F$ be an elliptic curve, $m\geq 2$, $k\geq 1$ and suppose that $p\nmid m\Delta_{F}$. If $F(\zeta_{p^{k}})\subseteq F(E[m])$, then we are in one of the following cases:
  • $k=1$ and $E/F$ has bad reduction at every ideal above $p$,
  • $k=2$, $p=2$ and at each prime above $p$, $E/F$ has either additive or non split multiplicative reduction,
  • $k=2$, $p=3$, and $E/F$ has additive and potential good reduction at every ideal above $p$,
  • $k=3$ or $4$, $p=2$ and $E/F$ has additive and potential good reduction at every ideal above $p$.
Proof.
Let $\mathfrak{p}$ be a prime ideal above $p$. Since $p\nmid\Delta_{F}$, the extension $F(\zeta_{p^{k}})/F$ is ramified at $\mathfrak{p}$, from Lemma 2.5, so is $F(E[m])/F$ by assumptions. Looking at the tables in Theorem 2.6 and Remark 2.7, with $e=1$, we see that the possibilities are: $k=1$, corresponding to the second line of each tables; $k=2$, corresponding to the third line and in this case $p=2,3$; or $k=3,4$, corresponding to the fourth line, where $p=2$. ∎
The corollary below tells us that if $E/F$ has a $(m,n)$-coincidence (with some conditions on $F$), then the primes greater than $5$ not dividing $m$ (respectively $n$) divide $n$ (respectively $m$) to at most power $1$, and $E/F$ must have bad reduction at these primes. Moreover, if $3\nmid m$, then $3$ divides $n$ to at most power $2$, and if $m$ is odd, then $2$ divides $n$ to at most power $4$, and the greater the power, the more restrictive is the reduction type.
Corollary 2.9.
Let $E/F$ be an elliptic curve with an $(m,n)$-coincidence. Suppose that $p\mid n$ and $p\nmid m\Delta_{F}$. Then we are in one of the following cases:
  • $v_{p}(n)=1$ and $E/F$ has bad reduction at every ideal above $p$,
  • $v_{p}(n)=2$, $p=2$ and at each prime above $p$, $E/F$ has either additive or non split multiplicative reduction,
  • $v_{p}(n)=2$, $p=3$, and $E/F$ has additive and potential good reduction at every ideal above $p$,
  • $v_{p}(n)=3$ or $4$, $p=2$ and $E/F$ has additive and potential good reduction at every ideal above $p$.
Remark 2.10.
If $e_{\mathfrak{p}}(F/\mathbb{Q})$ is prime to $\varphi(p^{k})$ (hypothesis that is satisfied for example if $F/\mathbb{Q}$ is unramified at $\mathfrak{p}$), then Corollaries 2.8 and 2.9 are true replacing "at every ideal above $p$" by "at $\mathfrak{p}$".
3. Coincidences in towers
In this section, we deal with coincidence in towers, or vertical coincidences, that is to say $(p^{k},p^{k+1})$-coincidences for a prime $p$ and a positive integer $k$. More generally, the section also contains results about $(m,n)$-coincidence where $m\mid n$.
3.1. Construction of vertical coincidences
Over $\mathbb{Q}$, we know that infinitely many elliptic curves have a $(2,4)$-coincidence and this is the only vertical coincidence which occurs, see [DLR23, Theorem 1.4]. Over a number field, there are additional possibilities. Obviously, to obtain an $(m,mn)$-coincidence for an elliptic curve $E/F$, it suffices to do a base change of the ground field to $F(E[mn])$. However, such a base change is a trivial construction and so not very relevant. We will say that the base change from $F$ to $L$ is minimal for an $(m,mn)$-coincidence if $L(E[m])=F(E[mn])$ and $F(E[m])\cap L=F$. Here is an example of a $(4,8)$-coincidence obtained by a minimal base change:
Theorem 3.1.
There are infinitely many isomorphism classes of elliptic curves $E/\mathbb{Q}$ such that there exists a number field $L$ with Galois group $(\mathbb{Z}/2\mathbb{Z})^{r}$ over $\mathbb{Q}$ with $1\leq r\leq 4$ satisfying
$L(E[4])=L(E[8])\neq L.$
Proof.
We apply Proposition 3.4 for $F=\mathbb{Q}$, $m=4$, $r=2$ and $E/\mathbb{Q}$ such that $\operatorname{\mathrm{Gal}}(\mathbb{Q}(E[8])/\mathbb{Q})\simeq(\mathbb{Z}/2\mathbb{Z})^{t}$ for some $t$. Hence, there exists $L/\mathbb{Q}$ of degree dividing #$\#\mathrm{GL}_{2}(8)/\#\mathrm{GL}_{2}(4)=2^{4}$ (by (3) in Subsection 3.3) such that $L(E[8])=L(E[4])\neq L$. By [GLR16, Theorem 1.1], the Galois group $\operatorname{\mathrm{Gal}}(\mathbb{Q}(E[8])/\mathbb{Q})$ is isomorphic to $(\mathbb{Z}/2\mathbb{Z})^{t}$ for some $t$ for infinitely many isomorphism classes of elliptic curves $E/\mathbb{Q}$ and $\mathbb{Q}(E[4])/\mathbb{Q}$ is non trivial since it contains $\zeta_{4}$, which completes the proof. ∎
Remark 3.2.
The proof of Theorem 3.1 considers only elliptic curves $E/\mathbb{Q}$ such that $\mathbb{Q}(E[8])/\mathbb{Q}$ is abelian. In this case $\operatorname{\mathrm{Gal}}(\mathbb{Q}(E[8])/\mathbb{Q})\simeq(\mathbb{Z}/2\mathbb{Z})^{t}$ with $t\in\{4,5,6\}$ from [GLR16, Theorem 1.1]. Let $r$ be as in Theorem 3.1. We have $2^{r}=\#\operatorname{\mathrm{Gal}}(\mathbb{Q}(E[8])/\mathbb{Q}(E[4]))$ by construction of $F$ and $1\leq r\leq 4$. If $t=4$, then $1\leq r\leq 3$ and if $t=6$ then $2\leq r\leq 4$.
Remark 3.3.
We cannot use the abelian case to construct $(p,p^{2})$-coincidence with $p$ odd, because there is no abelian $p^{2}$-division field for $p$ odd and $E$ defined over $\mathbb{Q}$.
Proposition 3.4.
Let $m,n$ be positive integers. Let $E/F$ be an elliptic curve such that the following exact sequence is split:
$1\to\operatorname{\mathrm{Gal}}(F(E[mn]/F(E[m]))\to\operatorname{\mathrm{Gal}}(F(E[mn]/F)\to\operatorname{\mathrm{Gal}}(F(E[m]/F)\to 1$
with $F(E[m])/F$ non trivial. Then there exists an extension $L/F$ of degree dividing #$\#\mathrm{GL}_{2}(mn)/\#\mathrm{GL}_{2}(m)$ such that
$L(E[m])=L(E[mn])\neq L.$
To prove the proposition, we will use the following elementary remark:
Remark 3.5.
Let $G$ be a group and $H$ be a subgroup of $G$ of finite index. Let $\phi:G\to G^{\prime}$ be a surjective morphism (of groups) and set $H^{\prime}=\phi(H)$. Then $\phi$ induces a surjective morphism of $G$-sets $G/H\to G^{\prime}/H^{\prime}$, from which
$[G:H]=[G^{\prime}:H^{\prime}][\ker(\phi)H:\ker(\phi)].$
In particular $[G^{\prime}:H^{\prime}]$ divides $[G:H]$.
Proof of Proposition 3.4.
Since the sequence is split, there exists a morphism
$\iota:\operatorname{\mathrm{Gal}}(F(E[m])/F)\to\operatorname{\mathrm{Gal}}(F(E[mn])/F)$
such that the composition with the restriction map
$\operatorname{\mathrm{Gal}}(F(E[mn])/F)\to\operatorname{\mathrm{Gal}}(F(E[m])/F)$
is the identity. Let $L$ be the fixed field of $\mathrm{Im}(\iota)$. Then $\operatorname{\mathrm{Gal}}(F(E[mn])/F)$ is the semi-direct product of $\operatorname{\mathrm{Gal}}(F(E[mn])/F(E[m]))$ by $\operatorname{\mathrm{Gal}}(F(E[mn])/L)$ and so $L(E[m])=F(E[mn])=L(E[mn])$ and $L\cap F(E[m])=F$. Since $F(E[m])/F$ is nontrivial, then $L(E[m])\neq L$. Moreover, the extension $L/F$ has degree $[F(E[mn]):F(E[m])]$, which divides #$\#\mathrm{GL}_{2}(mn)/\#\mathrm{GL}_{2}(m)$ by point (1) of Remark 3.5, taking for $\phi$ the natural map $\mathrm{GL}_{2}(mn)\to\mathrm{GL}_{2}(m)$ and $H=\rho_{E,mn}(G_{F})$. ∎
Corollary 3.6.
For $E/F$ an elliptic curve, the following are equivalent:
  1. (1)
    The following sequence is split
    $1\to\operatorname{\mathrm{Gal}}(F(E[mn]/F(E[m]))\to\operatorname{\mathrm{Gal}}(F(E[mn]/F)\to\operatorname{\mathrm{Gal}}(F(E[m]/F)\to 1$
  2. (2)
    There exists an injective morphism
    $\iota:\operatorname{\mathrm{Gal}}(F(E[m])/F)\to\operatorname{\mathrm{Gal}}(F(E[mn])/F)$
    such that
    $\operatorname{\mathrm{Gal}}(F(E[mn])/F)=\iota(\operatorname{\mathrm{Gal}}(F(E[m])/F))\ltimes\operatorname{\mathrm{Gal}}(F(E[mn])/F(E[m])).$
  3. (3)
    There exists a minimal base change $L/F$ such that $E/L$ has an $(m,mn)$-coincidence.
In this case, $\iota(\operatorname{\mathrm{Gal}}(F(E[m])/F)=\operatorname{\mathrm{Gal}}(F(E[mn])/L)$.
Proof.
The equivalence between point $(1)$ and $(2)$ is immediate from the definitions of split exact sequence and semi-direct product. We know that $(1)\implies(3)$ by Proposition 3.4. It remains to show that $(3)\implies(1)$. Suppose that the conditions of $(3)$ are satisfied. Since $F(E[mn])=L(E[mn])$, we have the following commutative diagram, where the horizontal arrows are restriction morphisms and the vertical arrows are inclusion morphisms:
${\operatorname{\mathrm{Gal}}(F(E[mn])/L)}$${\operatorname{\mathrm{Gal}}(L(E[m])/L)}$${\operatorname{\mathrm{Gal}}(F(E[mn])/F)}$${\operatorname{\mathrm{Gal}}(F(E[m])/F).}$$\scriptstyle{\psi}$$\scriptstyle{\phi}$
By assumption, $\psi$ is a isomorphism, together with $\phi$ by linear independance of $F(E[m])$ and $L$ over $F$. It follows that the exact sequence of $(1)$ splits by the morphism
$\begin{array}[]{ccccc}\iota&:&\operatorname{\mathrm{Gal}}(F(E[m])/F)&\longrightarrow&\operatorname{\mathrm{Gal}}(F(E[mn])/F)\\ &&\sigma&\longmapsto&(\phi\circ\psi)^{-1}(\sigma).\end{array}$
As a consequence of the corollary, the elliptic curves satisfying Theorem 3.1 are exactly those such that we have a split exact sequence
$1\to(\mathbb{Z}/2\mathbb{Z})^{r}\to\operatorname{\mathrm{Gal}}(\mathbb{Q}(E[8]/\mathbb{Q})\to\operatorname{\mathrm{Gal}}(\mathbb{Q}(E[4]/\mathbb{Q})\to 1$
In particular, this is true for $E/\mathbb{Q}$ such that $\mathbb{Q}(E[8])/\mathbb{Q}$ is a $(\mathbb{Z}/2\mathbb{Z})^{t}$-extension and a classification for such elliptic curves is given in [GLR16, Table 4]. But there are many other possibilities. More generally, we have
$\operatorname{\mathrm{Gal}}(F(E[p^{k+1}]/F(E[p^{k}]))\simeq(\mathbb{Z}/p\mathbb{Z})^{r}$
for some $r\leq 4$. Indeed, the Galois group $\operatorname{\mathrm{Gal}}(F(E[p^{k+1}]/F(E[p^{k}]))$ is isomorphic, for $n=2$, to a subgroup of
(2) $\ker\left(\mathrm{GL}_{n}(p^{k+1})\to\mathrm{GL}_{n}(p^{k})\right)=I_{n}+p^{k}\mathrm{M}_{n}(\mathbb{Z}/p^{k+1}\mathbb{Z})\simeq(\mathbb{Z}/p\mathbb{Z})^{n^{2}}.$
Hence, to construct a $(p^{k},p^{k+1})$-coincidence by minimal base change, we have and it suffices to find elliptic curves $E/F$ such that the following exact sequence is split:
$1\to(\mathbb{Z}/p\mathbb{Z})^{r}\to\operatorname{\mathrm{Gal}}(F(E[p^{k+1}]/F)\to\operatorname{\mathrm{Gal}}(F(E[p^{k}]/F)\to 1$
3.2. Trivial intersection with the cyclotomic field
In this section, we show that, if $F\cap\mathbb{Q}(\zeta_{p^{k}})$ is trivial, then a $(p^{k},p^{k+1})$-coincidence with $p$ prime is possible only for $p=2$.
Lemma 3.7.
Let $E/F$ be an elliptic curve, and $L/F$ be a cyclic extension such that $L\subseteq F(E[m])$. Let $\sigma\in G_{F}$ such that its restriction to $L$ generates $\operatorname{\mathrm{Gal}}(L/F)$. Then the order of $\rho_{E,m}(\sigma)$ is divisible by $[L:F]$.
Proof.
Let $\overline{\rho_{E,m}}$ be the reduction of $\rho_{E,m}$ modulo $\operatorname{\mathrm{Gal}}(\overline{F}/F(E[m]))$. Then
[…]
The first equality is by assumption, and the second is because of the injectivity of $\overline{\rho_{E,m}}$.∎
Theorem 3.8.
Let $E/F$ be an elliptic curve, $p$ be a prime and $k$ be a positive integer such that $F\cap\mathbb{Q}(\zeta_{p^{k}})=\mathbb{Q}$. If $F(\zeta_{p^{k+1}})\subseteq F(E[p^{k}])$, then $p=2$.
Proof.
Suppose that $p$ is odd and $F(\zeta_{p^{k+1}})\subseteq F(E[p^{k}])$. Let $\sigma\in G_{F}$ such that its restriction to $F(\zeta_{p^{k+1}})$ generates $\mathrm{Gal}(F(\zeta_{p^{k+1}})/F)$. Then its restriction to $F(\zeta_{p^{k}})$ generates $\operatorname{\mathrm{Gal}}(F(\zeta_{p^{k}})/F)$. So $\det\rho_{E,p^{k}}(\sigma)$ generates $(\mathbb{Z}/p^{k}\mathbb{Z})^{*}$. Moreover, Lemma 3.7 says that $\varphi(p^{k+1})$ divides the order of $\rho_{E,p^{k}}(\sigma)$ and so its determinant is a square mod $p$, by [DLR23, Lemma 3.5]. But, for $p$ odd, a square mod $p$ cannot generate $(\mathbb{Z}/p^{k}\mathbb{Z})^{*}$. Hence, $p$ is even. ∎
Remark 3.9.
The CM elliptic curve $y^{2}=x^{3}-11x-14$ satisfies $\mathbb{Q}(\zeta_{2^{k+1}})\subseteq\mathbb{Q}(E[2^{k}])$ for all $k\geq 1$. See [DLR23, Theorem 1.5], and [Jon23, Theorem 1.1] for more examples of such curves.
Corollary 3.10.
Let $E/F$ be an elliptic curve with $F\cap\mathbb{Q}(\zeta_{p^{k}})=\mathbb{Q}$. If $F(E[p^{k}])=F(E[p^{k+1}])$, then $p=2$.
Proof.
It is immediate from Theorem 3.8, since $F(\zeta_{p^{k+1}})\subseteq F(E[p^{k+1}])$. ∎
Remark 3.11.
In [DLR23, Theorem 1.4], Daniels and Lozano-Robledo have shown that, for $F=\mathbb{Q}$, only $k=1$ occurs.
Remark 3.12.
If $E/F$ has an $(m,mn)$-coincidence, then we must have $F(\zeta_{mn})\subseteq F(E[m])$. But this does not implies in general that $F(\zeta_{mn})=F(\zeta_{m})$, as we will see in Remark 3.24. Even more, unless $m$ is odd and $n=2$, this last never happens if $F=\mathbb{Q}$, and yet some coincidences occurs, like $(2,4)$ and $(2,6)$-coincidence, see [DLR23, Examples 1.2 and 1.3]. As in Remark 3.24, it is due to the non-surjectivity of $\mathrm{SL}_{2}(mn)\cap\rho_{E,mn}(G_{F})\to\mathrm{SL}_{2}(m)\cap\rho_{E,m}(G_{F})$.
Remark 3.13.
The condition $F(\zeta_{p^{k+1}})\subseteq F(E[p^{k}])$ is not sufficient to have the coincidence. For example, the elliptic curve of Remark 3.9 does not satisfy $\mathbb{Q}(E[2^{k}])=\mathbb{Q}(E[2^{k+1}])$ for any $k$. Indeed, this elliptic curve has CM and [DLR23, Proposition 3.9] implies that no CM elliptic curve defined over $\mathbb{Q}$ has a $(2^{k},2^{k+1})$-coincidence.
We are now able to prove the theorem stated in the introduction:
Theorem 3.14.
Let $m,n\geq 1$ and $E/F$ be an elliptic curve with conductor ideal $\mathfrak{f}_{E}$. Let $\mathrm{N}(\mathfrak{f}_{E})$ be the norm of $\mathfrak{f}_{E}$. Suppose that $F(E[m])=F(E[n])$. Then, for all primes $p$ such that $v_{p}(m)\neq v_{p}(n)$, we have
$p\mid 2\cdot\Delta_{F}\cdot\mathrm{N}(\mathfrak{f}_{E}).$
Proof.
First, suppose that $p$ divides $n$ or $m$ but not both. Then, by Corollary 2.9, if $p\nmid\Delta_{F}$, then $E/F$ has bad reduction above $p$. Now, suppose that $p$ divides both $n$ and $m$ such that $v_{p}(m)=k$ and $v_{p}(m)<v_{p}(n)$. Since $F(E[m])=F(E[\mathrm{lcm}(m,n)])$, then $F(E[m])=F(E[pm])$. Setting $a=\frac{m}{p^{k}}$ and $L=F(E[a])$, we obtain $F(E[p^{k}a])=F(E[p^{k+1}a])$ and $L(E[p^{k}])=L(E[p^{k+1}])$. Then Corollary 3.10 implies that $L\cap\mathbb{Q}(\zeta_{p^{k+1}})\neq\mathbb{Q}$ or $p=2$. In particular, $p$ is ramified in $L/\mathbb{Q}$ or $p=2$. But $L=F(E[a])$, so $p$ is ramified in $L/\mathbb{Q}$ if and only if $p$ is ramified in $F/\mathbb{Q}$ or in $F(E[a])/F$. Therefore $p\mid\Delta_{F}$ or $E$ has bad reduction above $p$. ∎
Corollary 3.15.
Let $m,n\geq 1$, $E/\mathbb{Q}$ be an elliptic curve and $\Delta_{E}$ be the minimal discriminant of $E$. Suppose that $\mathbb{Q}(E[m])=\mathbb{Q}(E[n])$. Then, for all primes $p$ such that $v_{p}(m)\neq v_{p}(n)$, we have $p\mid 2\Delta_{E}$.
3.3. Index of images
Let $p$ be a prime. Let $M$ be a subgroup of $\mathrm{GL}_{n}(\mathbb{Z}_{p})$ and let $G$ be a subgroup of $M$. In our setting we only need to consider the groups $\mathrm{SL}_{2}(\mathbb{Z}_{p})$, $\mathrm{GL}_{2}(\mathbb{Z}_{p})$ and $(\mathbb{Z}_{p})^{*}$, but we state results in the general case as the approach is the same. For $k$ a positive integer, we denote by $M_{k}$ the image of $M$ in $\mathrm{GL}_{n}(\mathbb{Z}/p^{k}\mathbb{Z})$ and $G_{k}$ the image of $G$ in $M_{k}$. We set $i_{k}=[M_{k}:G_{k}]$. We have $i_{k}\mid i_{k+1}$ by Remark 3.5. Moreover,
(3) $\frac{i_{k+1}}{i_{k}}=\frac{\#M_{k+1}}{\#G_{k+1}}\cdot\frac{\#G_{k}}{\#M_{k}}\left|\frac{\#M_{k+1}}{\#M_{k}}\right.=\left\{\begin{array}[]{ll}p^{n^{2}}&\mbox{if }M=\mathrm{GL}_{n}(\mathbb{Z}_{p})\\ p^{n^{2}-1}&\mbox{if }M=\mathrm{SL}_{n}(\mathbb{Z}_{p}).\end{array}\right.$
from (2). In particular,
(4) $G_{k}\simeq G_{k+1}\iff\frac{i_{k+1}}{i_{k}}=\left\{\begin{array}[]{ll}p^{n^{2}}&\mbox{if }M=\mathrm{GL}_{n}(\mathbb{Z}_{p})\\ p^{n^{2}-1}&\mbox{if }M=\mathrm{SL}_{n}(\mathbb{Z}_{p}).\end{array}\right.$
Proposition 3.16.
Let $E/F$ be an elliptic curve without CM, with a $(p^{k},p^{k+1})$-coincidence. Then $p^{4}$ divides $[\mathrm{GL}_{2}(\hat{\mathbb{Z}}):\rho_{E}(G_{F})]$.
Proof.
With the introduced notation, we consider $M=\mathrm{GL}_{2}(\mathbb{Z}_{p})$ and $G=\rho_{E,p^{\infty}}(G_{F})$. The index $i_{k+1}$ divides $[\mathrm{GL}_{2}(\hat{\mathbb{Z}}):\rho_{E}(G_{F})]$ by Remark 3.5 and so $i_{k+1}/i_{k}$ divides $[\mathrm{GL}_{2}(\hat{\mathbb{Z}}):\rho_{E}(G_{F})]$. But having a $(p^{k},p^{k+1})$-coincidence is equivalent to have $G_{k}\simeq G_{k+1}$. Then, the proposition follows from Equation 4. ∎
Remark 3.17.
In [Zyw22, Theorem 1.3], Zywina gives a set of possible index $[\mathrm{GL}_{2}(\hat{\mathbb{Z}}):\rho_{E}(G_{\mathbb{Q}})]$ which is generic for elliptic curves $E$ defined over $\mathbb{Q}$. In [Zyw24, Theorem 1.2], he gives upper bounds for the index $[\mathrm{GL}_{2}(\hat{\mathbb{Z}}):\rho_{E}(G_{F})]$ for elliptic curves $E/F$, depending on the number field $F$.
The idea of the lemma below and its proof follows [SZ17, Lemma 3.7].
Proposition 3.18.
Suppose that $M=\mathrm{GL}_{n}(\mathbb{Z}_{p})$ or $\mathrm{SL}_{n}(\mathbb{Z}_{p})$. The sequence $(u_{k})=\left(\frac{i_{k+1}}{i_{k}}\right)$ satisfies $u_{k+1}\mid u_{k}$ for $k\geq 1$ if $p$ is odd and for $k\geq 2$ if $p=2$.
Proof.
Suppose that $p$ is odd or that $k\geq 2$ and $p=2$. Let $H_{k}$ be the kernel of the reduction map $G_{k}\to G_{k-1}$. Let $h\in G$ whose image in $G_{k}$ belongs to $H_{k}$. Then $h=I+p^{k-1}A$ with $A\in M_{n}(\mathbb{Z}_{p})$. The map
$\phi:&H_{k}&\longrightarrow&H_{k+1}\\ &\overline{h}&\longmapsto&\overline{h^{p}}\end{array}$
is an injective morphism since
$(I+p^{k-1}A)^{p}=I+\binom{p}{1}p^{k-1}A+\binom{p}{2}p^{2{k-1}}A^{2}+\dots\equiv I+p^{k}A\ (\mathrm{mod}\ p^{k+1}).$
Therefore #$\frac{\#G_{k}}{\#G_{k-1}}$ divides #$\frac{\#G_{k+1}}{\#G_{k}}$ and so, since #$\frac{\#M_{k}}{\#M_{k-1}}=\frac{\#M_{k+1}}{\#M_{k}}$ from the equation (3), we obtain $u_{k+1}\mid u_{k}$. ∎
Corollary 3.19.
Let $k\geq 1$ if $p$ is odd and $k\geq 2$ if $p$ is even. If $M=\mathrm{GL}_{n}(\mathbb{Z}_{p})$ or $M=\mathrm{SL}_{n}(\mathbb{Z}_{p})$, and $G_{k}\simeq G_{k+1}$, then $G_{1}\simeq G_{2}\simeq\dots\simeq G_{k+1}$ if $p$ is odd, and $G_{2}\simeq G_{3}\simeq\dots\simeq G_{k+1}$ if $p$ is even.
Proof.
Suppose that $M=\mathrm{GL}_{n}(\mathbb{Z}_{p})$. Equivalence (3) gives $\frac{i_{k+1}}{i_{k}}=p^{n^{2}}$. Since the sequence $\left(\frac{i_{s+1}}{i_{s}}\right)$ is non-increasing from Lemma 3.18 for $s\geq 1$ and $p$ odd or $s\geq 2$ and $p=2$, and has values dividing $p^{n^{2}}$ by Equation (4), then $\frac{i_{s+1}}{i_{s}}=p^{4}$ for all $s\leq k$. The proof is similar for $M=\mathrm{SL}_{n}(\mathbb{Z}_{p})$. ∎
Theorem 3.20.
Let $q=p$ and $k\geq 1$ if $p$ is odd, or $q=p^{2}$ and $k\geq 2$ if $p$ is even. Let $E/F$ be an elliptic curve. If $F(E[p^{k}])=F(E[p^{k+1}])$, then $F(E[q])=F(E[p^{k+1}])$.
Proof.
Let $M=\mathrm{GL}_{2}(\mathbb{Z}_{p})$ and $G=\rho_{E,p^{\infty}}(G_{F})$. So
$G_{k}=\rho_{E,p^{k}}(G_{F})\simeq\operatorname{\mathrm{Gal}}(F(E[p^{k}])/F).$
Therefore, the equality $F(E[p^{k}])=F(E[p^{k+1}])$ is equivalent to $G_{k}\simeq G_{k+1}$, and we use Corollary 3.19. ∎
Remark 3.21.
Let $E/F$ be an elliptic curve and take $G=\rho_{E,p^{\infty}}(G_{F})$. So we have $G_{k}=\rho_{E,p^{k}}(G_{F})$ for $k\geq 1$.
  1. (1)
    The sequence $(i_{k})$ is increasing, and, if $E/F$ does not have CM, becomes stationary.
  2. (2)
    If $G=\mathrm{GL}_{2}(\mathbb{Z}_{p})$, then $\left(\frac{i_{k+1}}{i_{k}}\right)$ is constant equal to $1$. The converse is false: the elliptic curve with LMFDB label 11.a2 has non maximal Galois representation at $5$ and the sequence $\left(\frac{i_{k+1}}{i_{k}}\right)$ attached to $5$ is constant, equal to $1$.
  3. (3)
    For $E/\mathbb{Q}$ without CM, the $2$-adic image is the inverse image in $\mathrm{GL}_{2}(\mathbb{Z}_{2})$ of $\rho_{E,32}(G_{\mathbb{Q}})$, from [RZB15, Corollary 1.3]. Therefore, the sequence $\left(\frac{i_{k+1}}{i_{k}}\right)$ stabilizes at $1$ from $k$ at most $5$.
  4. (4)
    We consider $k\geq 1$ if $p$ is odd, or $k\geq 2$ if $p$ is even. If the first term of $\left(\frac{i_{k+1}}{i_{k}}\right)$ is different than $p^{4}$, then $E/F$ does not have $(p^{k},p^{k+1})$-coincidences for any $k$. Otherwise, if $s$ is the rank of the first jump of the sequence, then $E/F$ does not have $(p^{k},p^{k+1})$-coincidences for $k\geq s$.
  5. (5)
    We consider $k\geq 1$ if $p$ is odd, or $k\geq 2$ if $p$ is even. We know that the sequences $\left(\frac{i_{k+1}}{i_{k}}\right)$ is non-increasing, then constant. We can ask if it is decreasing then constant. The answer is no. For example, the elliptic curve with LMFDB label 15.a4 has, for $p=2$, $\left(\frac{i_{k+1}}{i_{k}}\right)=(2^{2},2,2,2,1,\dots)$. We can also ask if the graphs are "progressively non-increasing", meaning that $\frac{i_{k+1}}{i_{k}}\in\left\{\frac{i_{k}}{i_{k-1}},\frac{1}{p}\frac{i_{k}}{i_{k-1}}\right\}$. The answer is also no. For example, the elliptic curve with LMFDB label 15.a8 has, for $p=2$, $\left(\frac{i_{k+1}}{i_{k}}\right)=(2^{3},2,1,\dots)$ and the elliptic curve with LMFDB label 40.a4 has, for $p=2$, $\left(\frac{i_{k+1}}{i_{k}}\right)=(2^{4},1,\dots)$.
Example 3.22.
Here some examples over $\mathbb{Q}$, computed from [LMF24], illustrate different possibilities for the sequence $\left(\frac{i_{k+1}}{i_{k}}\right)$:
LMFDB Minimal Weierstrass equation Non max Sequence $\left(\frac{i_{k+1}}{i_{k}}\right)$
label $p$ attached to $p$
14.a6 $y^{2}+xy+y=x^{3}+4x-6$ $2$ $1,2,1,\dots$
15.a1 $y^{2}+xy+y=x^{3}+x^{2}-2160x-39540$ $2$ $2^{2},2^{2},2,1,\dots$
15.a2 $y^{2}+xy+y=x^{3}+x^{2}-135x-660$ $2$ $2^{2},2^{2},1,\dots$
15.a4 $y^{2}+xy+y=x^{3}+x^{2}-80x+242$ $2$ $2^{2},2,2,2,1,\dots$
15.a5 $y^{2}+xy+y=x^{3}+x^{2}-10x-10$ $2$ $2^{3},2,1,\dots$
15.a8 $y^{2}+xy+y=x^{3}+x^{2}+35x-28$ $2$ $2^{3},2^{2},1,\dots$
20.a3 $y^{2}=x^{3}+x^{2}-x$ $2$ $2,2,1,\dots$
40.a4 $y^{2}=x^{3}+13x-34$ $2$ $2^{4},1,\dots$
19.a1 $y^{2}+y=x^{3}+x^{2}-769x-8470$ $3$ $3,3,1,\dots$
54.a2 $y^{2}+xy=x^{3}-x^{2}-3x+3$ $3$ $3^{2},1,\dots$
11.a1 $y^{2}+y=x^{3}-x^{2}-7820x-263580$ $5$ $5,1,\dots$
11.a2 $y^{2}+y=x^{3}-x^{2}-10x-20$ $5$ $1,1,\dots$
Remark 3.23.
Let $E/F$ be an elliptic curve and take $G=\rho_{E,p^{\infty}}(G_{F})$. Then $\det G$ is a subgroup of $(\mathbb{Z}_{p})^{*}$, with image $\det G_{k}$ in $(\mathbb{Z}/p^{k}\mathbb{Z})^{*}$. We recall that $\det G_{k}\simeq\operatorname{\mathrm{Gal}}(F(\zeta_{p^{k}})/F)$, by Proposition 1.3. We set $j_{k}=[(\mathbb{Z}/p^{k}\mathbb{Z})^{*}:\det G_{k}]$. We consider $k\geq 1$ if $p$ is odd, and $k\geq 2$ is $p$ is even. The sequence $\left(\frac{j_{k+1}}{j_{k}}\right)$ is non-increasing and has value in $\{1,p\}$ by Lemma 3.18 and the equation (3). We have $j_{k+1}/j_{k}=p$ if and only if $\det G_{k}\simeq\det G_{k+1}$, by the isomorphism (4), and this is equivalent to $F(\zeta_{p^{k+1}})=F(\zeta_{p^{k}})$. Corollary 3.19 implies that $F(\zeta_{q})=F(\zeta_{p^{k+1}})$ with $q=p$ if $p$ is odd and $q=4$ if $p$ is even. As a consequence, the sequence $(j_{k})$ is increasing and becomes stationary from the smallest $s$ such that $\zeta_{p^{s}}\notin F(\zeta_{q})$.
Remark 3.24.
Let $E/F$ be an elliptic curve and take $G=\rho_{E,p^{\infty}}(G_{F})$. We observe that $\mathrm{SL}_{2}(\mathbb{Z}_{p})\cap G$ is a subgroup of $\mathrm{SL}_{2}(Z_{p})$ and so we can consider its projection in $\mathrm{SL}_{2}(\mathbb{Z}/p^{k}\mathbb{Z})$ for each $k$. Unfortunately, these projections are not necessarily equal to $\mathrm{SL}_{2}(p^{k+1})\cap G_{k}$ and so we cannot use these groups to deal with the coincidence : if $G_{k}\simeq G_{k+1}$, we do not necessarily have $\mathrm{SL}_{2}(p^{k})\cap G_{k}\simeq\mathrm{SL}_{2}(p^{k+1})\cap G_{k+1}$. Setting $\ell_{k}=[\mathrm{SL}_{2}(p^{k}):\mathrm{SL}_{2}(p^{k})\cap G_{k}]$, we have $i_{k}=j_{k}\ell_{k}$ with $j_{k}$ defined as in Remark 3.23. Suppose that $G_{k}\simeq G_{k+1}$. Then $i_{k+1}/i_{k}=p^{4}$. Hence
$p^{4}=\frac{j_{k+1}\ell_{k+1}}{j_{k}\ell_{k}}$ $=p^{4}\frac{\#\det G_{k}}{\#\det G_{k+1}}\frac{\#(\mathrm{SL}_{2}(p^{k})\cap G_{k})}{\#(\mathrm{SL}_{2}(p^{k+1})\cap G_{k+1})}$
$=p^{4}\frac{[F(\zeta_{p^{k}}):F]}{[F(\zeta_{p^{k+1}}):F]}\frac{\#(\mathrm{SL}_{2}(p^{k})\cap G_{k})}{\#(\mathrm{SL}_{2}(p^{k+1})\cap G_{k+1})}\cdot$
Then we have two situations:
  1. (1)
    $F(\zeta_{p^{k}})=F(\zeta_{p^{k+1}})$, and $\mathrm{SL}_{2}(p^{k})\cap G_{k}\simeq\mathrm{SL}_{2}(p^{k+1})\cap G_{k+1}$,
  2. (2)
    $F(\zeta_{p^{k+1}})\neq F(\zeta_{p^{k+1}})$, and the reduction map $\mathrm{SL}_{2}(p^{k+1})\cap G_{k+1}\to\mathrm{SL}_{2}(p^{k})\cap G_{k}$ is not surjective.
In the first case, we have seen in Remark 3.23 that $F(\zeta_{p^{k+1}})$ is equal to $F(\zeta_{4})$ if $p=2$ and $F(\zeta_{p})$ otherwise. The examples of vertical coincidence we have for elliptic curves over $\mathbb{Q}$ fits, obviously, in the second case. Indeed, the elliptic curve with LMFDB label 40.a4 has a $(2,4)$-coincidence with
$G_{2}=\rho_{E,4}(G_{F})\simeq\left\langle\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\right\rangle.$
In this case, we have $G_{2}\cap\mathrm{SL}_{2}(4)=\{\operatorname{\mathrm{id}}\}$, whereas
$G_{1}\cap\mathrm{SL}_{2}(2)=G_{1}=\left\langle\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\right\rangle.$
3.4. Split liftable subgroups
Let $m,n$ be positive integers. If $E/F$ is an elliptic curve with an $(m,mn)$-coincidence, then $\rho_{E,m}(G_{F})\simeq\rho_{E,mn}(G_{F})$ and the image of $\rho_{E,mn}(G_{F})$ in $\mathrm{GL}_{2}(m)$ is $\rho_{E,m}(G_{F})$. It leads to the following definition. For a subgroup $G$ of $\mathrm{GL}_{2}(m)$, there are a priori several liftings of $G$ in $\mathrm{GL}_{2}(mn)$.
Definition 3.25.
We say that a subgroup $G$ of $\mathrm{GL}_{2}(m)$ is split liftable modulo $mn$ if there exists $G^{\prime}\leq\mathrm{GL}_{2}(mn)$ such that $G$ is the image of $G^{\prime}$ in $\mathrm{GL}_{2}(m)$ and $G\simeq G^{\prime}$. We say that an element $g$ of $\mathrm{GL}_{2}(m)$ is split liftable modulo $m$ if there exists $g^{\prime}\in\mathrm{GL}_{2}(mn)$ with same order as $g$ and such that $g$ is the image of $g^{\prime}$ in $\mathrm{GL}_{2}(m)$.
A subgroup $G$ of $\mathrm{GL}_{2}(m)$ is split liftable modulo $mn$ is there exists an injective morphism $G\to\mathrm{GL}_{2}(mn)$ which makes the following diagram commutative:
${\mathrm{GL}_{2}(mn)}$${G}$${\mathrm{GL}_{2}(m)}$
The definition above is up to conjugation, since two conjugate groups are isomorphic. Therefore:
Proposition 3.26.
Let $E/F$ be an elliptic curve with an $(m,mn)$-coincidence. Then $\rho_{E,m}(G_{F})$ is split liftable modulo $mn$.
The aim of this section is to determine the subgroups of $\mathrm{GL}_{2}(m)$ which are split liftable or not modulo some multiple of $m$.
Remark 3.27.
If $G$ in $\mathrm{GL}_{2}(m)$ is split liftable modulo $mn$, then $G$ is split liftable modulo every $km$ such that $1\leq k\leq n$.
Remark 3.28.
In [Elk06], Elkies already use the property of being split liftable to construct the modular curve $\mathcal{X}_{9}$. It is defined by $\mathcal{X}_{9}=X(9)/(G/\left\langle-\operatorname{\mathrm{id}}\right\rangle)$ where $G$ is a split lifting of $\mathrm{SL}_{2}(3)$ in $\mathrm{GL}_{2}(9)$, and more specifically, in $\mathrm{SL}_{2}(9)$.
Remark 3.29.
Corollary 3.19 tells that, if a subgroup of $\mathrm{GL}_{2}(p^{k})$ is split liftable modulo $\mathrm{GL}_{2}(p^{k+1})$, then its image in $\mathrm{GL}_{2}(q)$ is also split liftable modulo $\mathrm{GL}_{2}(p^{k+1})$, where $q=p$ if $p$ is odd or $q=4$ if $p$ is even.
Lemma 3.30.
Let $G\leq\mathrm{GL}_{2}(m)$ be split liftable modulo $mn$. Then, all subgroups of $G$ are split liftable modulo $mn$.
Proof.
Let $G^{\prime}$ be a subgroup of $\mathrm{GL}_{2}(mn)$ such that $G$ is the image of $G^{\prime}$ in $\mathrm{GL}_{2}(mn)$ and $G\simeq G^{\prime}$. Then, the restriction $\pi:G^{\prime}\to G$ of the natural projection $\mathrm{GL}_{2}(mn)\to\mathrm{GL}_{2}(m)$ is an isomorphism. Let $H\leq G$. We set $H^{\prime}:=\pi^{-1}(H)$. Then $H^{\prime}\simeq H$ and $H$ is the image of $H^{\prime}$ in $\mathrm{GL}_{2}(m)$. ∎
Proposition 3.31.
Let $m\geq 2$. The following subgroups of $\mathrm{GL}_{2}(m)$ are split liftable modulo every multiple of $m$:
$\left\langle\begin{pmatrix}0&-1\\ 1&1\end{pmatrix},\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\right\rangle,\quad\left\langle\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},\begin{pmatrix}0&1\\ -1&0\end{pmatrix}\right\rangle.$
Proof.
Looking the first two groups as subgroups of $\mathrm{GL}_{2}(\mathbb{Z})$, we observe that they have finite orders, respectively $12$ and $8$, and their elements only have coefficients in $\{0,1,-1\}$. Consequently, it is isomorphic to their projection modulo any integers $m$ such that $1\neq-1\ (\mathrm{mod}\ m)$, that is for any $m\geq 3$. The case $m=2$ is given by Example 3.32. ∎
Example 3.32.
The group $\mathrm{GL}_{2}(2)$ lifts to $\left\langle\begin{pmatrix}-1&1\\ -1&0\end{pmatrix},\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\right\rangle\subseteq\mathrm{GL}_{2}(\mathbb{Z})$, and so is split liftable modulo every even integer.
Remark 3.33.
Let $E/F$ be an elliptic curve. To have an $(m,mn)$-coincidence, it is necessary to have $\rho_{E,m}(G_{F})$ split liftable modulo $mn$, but it is not sufficient. Indeed, $\mathrm{GL}_{2}(2)$ is split liftable modulo $8$, but there are no $(2,8)$-coincidence for elliptic curve defined over $\mathbb{Q}$, see [DLR23, Theorem 1.4]. This is also the case for $\mathrm{GL}_{2}(3)$: the subgroup
$\left\langle\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},\begin{pmatrix}-2&2\\ -2&-2\end{pmatrix},\begin{pmatrix}4&-2\\ -3&4\end{pmatrix}\right\rangle\subseteq\mathrm{GL}_{2}(9)$
is a lifting of $\mathrm{GL}_{2}(3)$ of order $48$. Hence, $\mathrm{GL}_{2}(3)$ is split liftable modulo $9$ and yet there is no $(3,9)$-coincidence for elliptic curves with surjective mod $3$ Galois representation, by Corollary 3.10 and Proposition 1.3.
Now we will present groups which are not split liftable, which give us obtructions having the coincidence. We underline that, by Lemma 3.30, if $g\in\mathrm{GL}_{2}(m)$ is not split liftable modulo $mn$, then all groups containing $g$ are not split liftable modulo $mn$.
From now on, we will denote by $T$ the matrix $\begin{pmatrix}1&1\\ 0&1\end{pmatrix}$.
Lemma 3.34.
Let $p$ be a prime and $k\geq 1$. The matrix $T$ in $\mathrm{GL}_{2}(p^{k})$ is split liftable modulo $p^{k+1}$ if and only if $p=2,3$ and $k=1$.
Proof.
In $\mathrm{GL}_{2}(2)$, resp. $\mathrm{GL}_{2}(3)$, the matrix $T$ is conjugates to $\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$, resp. $\begin{pmatrix}0&1\\ -1&-1\end{pmatrix}$, and so is split liftable modulo $4$, resp. modulo $9$, by Lemma 3.31. Set $q=p^{2}$ if $p=2,3$ and $q=p$ otherwise. We know, from Corollary 3.19, that, for $k\geq 1$ if $p$ is odd and $k\geq 2$ if $p$ is even, if $T$ in $\mathrm{GL}_{2}(p^{k})$ is split liftable modulo $p^{k+1}$, then $T$ in $\mathrm{GL}_{2}(q)$ is split liftable modulo $p^{k+1}$ and so modulo $pq$. Now, if $T$ was split liftable modulo $pq$, then we could find $M\in M_{2}(\mathbb{Z}_{p})$ such that $T+qM$ has order $q$ in $\mathrm{GL}_{2}(pq)$. But
$\left(\begin{smallmatrix}1+qa&1+qb\\ qc&1+qd\end{smallmatrix}\right)^{n}\equiv\left(\begin{smallmatrix}1+nqa+\frac{n(n-1)}{2}qc\quad&n+nqb+\frac{n(n-1)}{2}q(a+d)+\frac{n(n-1)(n-2)}{6}qc\\ nqc&1+nqd+\frac{n(n-1)}{2}qc\end{smallmatrix}\right)\ (\mathrm{mod}\ pq).$
Now, we take $n=q$. Then $p$ divides $\frac{q(q-1)}{2}$ and $\frac{q(q-1)(q-2)}{6}$. We obtain
$\left(T+qM\right)^{q}\equiv\begin{pmatrix}1&q\\ 0&1\end{pmatrix}\not\equiv\operatorname{\mathrm{id}}\ (\mathrm{mod}\ pq).$
Theorem 3.35.
Let $p$ be a prime. Let $q=p$ if $p\neq 2,3$, or $q=p^{2}$ if $p=2,3$. Let $E/F$ be an elliptic curve. If $\rho_{E,q}(G_{F})$ contains $T$, then $F(E[q])\neq F(E[pq])$.
In particular, if $E/F$ does not have CM and has a $(p^{k},p^{k+1})$ coincidence, then $E/F$ has non maximal image modulo $p^{2}$, and even modulo $p$ if $p\neq 2,3$.
Corollary 3.36.
Let $E/F$ be an elliptic curve with multiplicative reduction at a prime $\mathfrak{r}$ of $\mathcal{O}_{F}$ and let $p$ be a prime not dividing $2v_{\mathfrak{r}}(j(E))$. Then $E/F$ does not have a $(p^{k},p^{k+1})$-coincidence for any $k$.
Proof.
If $E/F$ has multiplicative reduction at a prime ideal $\mathfrak{r}$ and $p\nmid 2v_{\mathfrak{r}}(j(E))$, then $T\in\rho_{E,p}(G_{F})$ by [Sil94, Proposition 1.6]. But this is not possible for a $(p^{k},p^{k+1})$-coincidence from Theorem 3.35. ∎
Remark 3.37.
Suppose that $p$ is odd, and $k\geq 2$ if $p=3$. To study $(p^{k},p^{k+1})$-coincidences further, it remains to deal with subgroups of $\mathrm{GL}_{2}(p^{k})$ with non-surjective determinant, by Corollary 3.10, and which does not contains $T$, by Theorem 3.35.
3.5. CM case
If $E/F$ has complex multiplication by a quadratic field $K$ and $F\subseteq K(j(E))$, then we can say more.
Proposition 3.38.
Let $E/\overline{\mathbb{Q}}$ be an elliptic curve with CM by a quadratic field $K$ and $F=K(j(E))$. If $F(E[p^{k}])=F(E[p^{k+1}])$, then $p=2$ and $k=1$.
Proof.
We have the following field inclusions:
$F(E[p^{k+1}])$   $F(E[p^{k}])$$F(h(E[p^{k+1}])$$F(h(E[p^{k}]))$$a$$c$$d$
where $h$ is a Weber function for $E$ (see [LR22]). Suppose that $k\geq 1$ if $p$ is odd or $k\geq 2$ if $p=2$ and $a=1$. By [LR22, Theorem 4.3], we have,
$d=[F(h(E[p^{k+1}])):F(h(E[p^{k}]))]=p^{2}.$
This implies that $p^{2}\mid c$. Moreover, by [LR22, Theorem 4.1] we have $c\mid\#\mathcal{O}_{K}^{*}$. But #$\#\mathcal{O}_{K}^{*}=2$, $4$ or $6$. Thus $p=2$ and $c=\#\mathcal{O}_{K}^{*}=4$. But $\mathcal{O}_{K}^{*}\simeq\operatorname{\mathrm{Aut}}(E)$, so $j(E)=1728$ by [Sil09, III, Theorem 10.1] and $F=K(j(E))=K=\mathbb{Q}(i)$. In this case, $E$ is defined over $\mathbb{Q}$ and $\mathbb{Q}(E[2^{k}])\subsetneq\mathbb{Q}(E[2^{k+1}])$ by [DLR23, Proposition 3.9]. Moreover, the Weil pairing implies that
$F=\mathbb{Q}(i)\subseteq\mathbb{Q}(E[4])\subseteq\mathbb{Q}(E[2^{k}])\subsetneq\mathbb{Q}(E[2^{k+1}])$
and so $F(E[2^{k}])\subsetneq F(E[2^{k+1}])$. We conclude that $p=2$ and $k=1$. ∎
Remark 3.39.
If an elliptic curve $E/F$ has a $(2,4)$-coincidence, then $\rho_{E,4}(G_{F})$ must be a split lifting of $\rho_{E,2}(G_{F})$. Example 3.32 gives such a split lifting and a Magma computation shows that this is the only one up to conjugation ([DLR23, Proof of Proposition 3.9]). The corresponding modular curve is $X_{20b}$ in the notation of Rouse and Zureick-Brown [RZB15, Remark 1.6]. They have computed its model, see https://users.wfu.edu/rouseja/2adic/X20b.html, and so the map to the $j$-line, explicitely given in [DLR23, Proof of Proposition 3.9]. If the elliptic curve $E/F$ with $j$-invariant $j(E)$ has a $(2,4)$-coincidence, then there exists $t\in F$ such that
$j(E)=\frac{-4t^{8}+32t^{7}+80t^{6}-288t^{5}-504t^{4}+864t^{3}+1296t^{2}-864t-1188}{t^{4}+4t^{3}+6t^{2}+4t+1}\cdot$
For rational CM $j$-invariant, there is no such $t$.
4. Large images
Since $\mathrm{GL}_{2}(m)$ and $\mathrm{GL}_{2}(n)$ are not isomorphic for $m\neq n$, then $(m,n)$-coincidences cannot happen for elliptic curves with surjective mod $m$ and mod $n$ representations. In this section, we show that, under some conditions on $F$, $(m,n)$-coincidences cannot happen if one of the images, said the image mod $m$, is large, i.e. it contains the special linear group. In this case, the elliptic curve $E/F$ does not have CM, and it is said that it has maximal image at $m$. We will only deal with $m$ odd.
For a group $G$, we denote by $\operatorname{\mathrm{D}}(G)$ its commutator subgroup. We know that $D(G)$ is the smallest normal subgroup of $G$ such that $G/\operatorname{\mathrm{D}}(G)$ is abelian and this last is called the abelianization of $G$. We will use several well-known results about derived group of $\mathrm{SL}_{2}(m)$ and $\mathrm{GL}_{2}(m)$, given with detailed proofs in Appendix A.
We denote by $F^{\mathrm{ab}}$ the maximal abelian extension of $F$. We will compare the maximal abelian extension of $F(E[m])$ and that of $F(E[n])$.
Proposition 4.1.
Let $m$ be an odd integer and $E/F$ be an elliptic curve. Suppose that $\rho_{E,m}(G_{F})$ contains $\mathrm{SL}_{2}(m)$. Then
[…]
Proof.
Let $G=\rho_{E,m}(G_{F})$. We have $\mathrm{SL}_{2}(m)\leq G\leq\mathrm{GL}_{2}(m)$. Then, using Proposition A.4,
$\operatorname{\mathrm{D}}(\mathrm{SL}_{2}(m))\leq\operatorname{\mathrm{D}}(G)\leq\mathrm{SL}_{2}(m)).$
Suppose that $\operatorname{\mathrm{D}}(G)=\mathrm{SL}_{2}(m)$. Since
$G/\mathrm{SL}_{2}(m)\simeq\det(G),$
therefore the largest abelian quotient of $F(E[m])/F$ has Galois group isomorphic to $\det(G)$. By the Weil pairing, $F(\zeta_{m})\subseteq F(E[m])$ and from Proposition 1.3, we have $\operatorname{\mathrm{Gal}}(F(\zeta_{m})/F)\simeq\det(G)$. Then the largest abelian subextension of $F(E[m])$ is $F(\zeta_{m})$.
Now, suppose that $\operatorname{\mathrm{D}}(G)\neq\mathrm{SL}_{2}(m)$. If $3\nmid m$, this does not happens, since in this case $\operatorname{\mathrm{D}}(\mathrm{SL}_{2}(m))=\mathrm{SL}_{2}(m)$ from Proposition A.2. If $3\mid m$, then $\operatorname{\mathrm{D}}(\mathrm{SL}_{2}(m))$ has index $3$ in $\mathrm{SL}_{2}(m)$ from Proposition A.2 and so is $\operatorname{\mathrm{D}}(G)$. It follows that $F(E[m])\cap F^{\mathrm{ab}}$ is an extension of degree $3$ of $F(\zeta_{m})$. ∎
Remark 4.2.
The case $F(E[m])\cap F^{\mathrm{ab}}\neq F(\zeta_{m})$ happens only for $\gcd(m,12)\neq 1$ by Proposition A.2 and Proposition A.4. Let $k=v_{3}(m)$. In the previous proposition, $m$ is odd and so $F(E[m])\cap F^{\mathrm{ab}}\neq F(\zeta_{m})$ only if $k>0$. In this case $L:=F(E[3^{k}])\cap F^{\mathrm{ab}}$ is a $(\mathbb{Z}/3\mathbb{Z})$-extension of $F(\zeta_{3^{k}})$ and
$F(E[m])\cap F^{\mathrm{ab}}=L\otimes_{F}F(\zeta_{\frac{m}{3^{k}}}).$
Remark 4.3.
For $p\geq 5$, if $\rho_{E,p}(G_{F})$ contains $\mathrm{SL}_{2}(p)$, then $\rho_{E,p^{\infty}}(G_{F})$ contains $\mathrm{SL}_{2}(\mathbb{Z}_{p})$, by [Ser89, IV.3.4.Lemma 3].
Theorem 4.4.
Let $m$ be an odd integer, $n\nmid m$ such that $\zeta_{n}\notin F$ and $E/F$ be an elliptic curve. Suppose that $\rho_{E,m}(G_{F})$ contains $\mathrm{SL}_{2}(m)$ and that $E/F$ has an $(m,n)$-coincidence. Then
$3\mid m,\quad\text{and}\quad\mathrm{D}(\rho_{E,m}(G_{F}))\neq\mathrm{SL}_{2}(m),\quad\text{and}\quad F(\zeta_{n})\subseteq L$
with $L$ a $\mathbb{Z}/3\mathbb{Z}$-extension of $F(\zeta_{m})$.
Proof.
Suppose that $F(E[m])=F(E[n])$. Then $F(\zeta_{n})\subseteq F(E[m])\cap F^{\mathrm{ab}}$. If $F(E[m])\cap F^{\mathrm{ab}}=F(\zeta_{m})$, this is not possible since $\zeta_{n}\notin F$ and $n\nmid m$. Hence, from Proposition 4.1, $F(E[m])\cap F^{\mathrm{ab}}$ is a $\mathbb{Z}/3\mathbb{Z}$-extension of $F(\zeta_{m})$ and the derived group of $\rho_{E,m}(G_{F})$ is smaller than $\mathrm{SL}_{2}(m)$, which only happen if $3\mid m$ by Remark 4.2. ∎
Remark 4.5.
Under the hypotheses of the previous theorem, for $(m,n)=(p^{k},p^{k+1})$, we know by the previous theorem that $p=2$ or $3$. Since $\mathrm{SL}_{2}(p^{k})$ contains the matrix $T$, this results was already known by Theorem 3.35 and even more: the assumption $\zeta_{p^{k+1}}\notin F$ is unnecessary and $k=1$.
We finish the section by the following lemma, which contains some additional information about the extension $F(E[3])/F$.
Lemma 4.6.
Let $E/F$ be an elliptic curve with $j$-invariant $j(E)$. We have $F(j(E)^{1/3})\subseteq F(E[3])$. Moreover, if $\mathrm{SL}_{2}(3)\subseteq\rho_{E,3}(G_{F})$, then $F(j(E)^{1/3})/F$ is non trivial.
Proof.
We know that we can parameterize the modular curve $X(1)$ with $j:X(1)\to\mathbb{P}^{1}$ the $j$-invariant. Let $C_{ns}^{+}(3)$ be the normalizer of the non-split Cartan subgroup of $\mathrm{GL}_{2}(3)$ and the associate modular curve $X_{ns}^{+}(3)=X(3)/C_{ns}^{+}(3)$. From [Che99, Proposition 4.1], there exists a uniformizer $t:X_{ns}^{+}(3)\to\mathbb{P}^{1}$ such that $t^{3}=j$. The points of $X_{ns}^{+}(3)(F(E[3]))$ correspond to elliptic curves defined over $F(E[3])$ with Galois image contained in $C_{ns}^{+}(3)$, that is every elliptic curves defined over $F$, since $\rho_{E,3}(G_{F(E[3])})=\operatorname{\mathrm{id}}$. Hence $t(E)\in F(E[3])$. Using that $t^{3}=j$, we obtain $j(E)^{1/3}\in F(E[3])$. Moreover, if $j(E)^{\frac{1}{3}}\in F$, then $\rho_{E,3}(G_{F})\subseteq C_{ns}^{+}(3)$. In particular, in this case, $\rho_{E,3}(G_{F})$ cannot contain $\mathrm{SL}_{2}(3)$. ∎
Remark 4.7.
If $\zeta_{3}\in F$, then $F(j(E)^{\frac{1}{3}})/F$ is Galois and so is contained in $F^{\mathrm{ab}}$. If, moreover, $F(j(E)^{\frac{1}{3}})\cap F(\zeta_{m})=F$, then $F(E[m])\cap F^{\mathrm{ab}}=F(j(E)^{\frac{1}{3}},\zeta_{m})$. Hence, if we have an $(m,n)$-coincidence, then
$3\mid m,\quad\text{and}\quad\mathrm{D}(\rho_{E,m}(G_{F}))\neq\mathrm{SL}_{2}(m),\quad\text{and}\quad F(\zeta_{n})\subseteq F(j(E)^{\frac{1}{3}},\zeta_{m}).$
Remark 4.8.
If $\zeta_{3}\in F$, $3\mid m$ and $\operatorname{\mathrm{D}}(\rho_{E,3^{v_{3}(m)}}(G_{F}))=\mathrm{SL}_{2}(3^{v_{3}(m)})$, then $F(j(E)^{\frac{1}{3}})\subseteq F(\zeta_{3^{v_{3}(m)}})$ and so we have $r\geq 1$ such that
$F(j(E)^{\frac{1}{3}})=F(\zeta_{3^{r+1}})\quad\text{and}\quad F=F(\zeta_{3^{r}})\neq F(\zeta_{3^{r+1}}).$
5. Coincidence of division fields of two elliptic curves
Let $E$ and $E^{\prime}$ be elliptic curves defined over $F$. In all this section, we suppose that $F(E[m])=F(E^{\prime}[n])$ for two integers $m$ and $n$. Then $\zeta_{n}\in F(E[m])$ and in the previous sections we gave constraints to this property. Hence:
Proposition 5.1.
Let $p$ be a prime and $r$ be the largest integer such that $\mathbb{Q}(\zeta_{p^{r}})\subseteq F\cap\mathbb{Q}(\zeta_{p^{\infty}})$. Suppose that $p>q$ for all primes $q\mid m$ and $v_{p}(n)>r$. Then, $v_{p}(n)=1$ (and $r=0$), unless $(m,p)=(2^{j},3)$ for some $j\geq 1$, in which case $r=0$ and $v_{p}(n)\leq 2$, or $v_{p}(n)=r+1$.
Proof.
This follows from Proposition 1.5 and the fact that $F(\zeta_{p^{k}})\subseteq F(E^{\prime}[p^{k}])\subseteq F(E^{\prime}[n])$. ∎
We can use Theorem 2.6 and Remark 2.7 as the same way as in Section 2 to give constraints on a coincidence $F(E[m])=F(E^{\prime}[n])$ using again that we must have $\zeta_{n}\in F(E[m])$. For example, we give a generalization of Corollary 2.9.
Theorem 5.2.
For all primes $p$ such that $p\mid n$ and $p\nmid m\Delta_{F}$ we are in one of the following situation:
  • $v_{p}(n)=1$ and $E/F$ has bad reduction at every ideal above $p$,
  • $v_{p}(n)=2$, $p=2$ and at each prime above $p$, $E/F$ has either additive or non split multiplicative reduction,
  • $v_{p}(n)=2$, $p=3$, and $E/F$ has additive and potential good reduction at every ideal above $p$,
  • $v_{p}(n)=3$ or $4$, $p=2$ and $E/F$ has additive and potential good reduction at every ideal above $p$.
Theorem 5.3.
Let $p$ be a prime and $k\geq 1$ such that $F\cap\mathbb{Q}(\zeta_{p^{k}})=\mathbb{Q}$. If $F(E[p^{k}])=F(E^{\prime}[p^{k+1}])$, then $p=2$.
Proof.
This follows from Theorem 3.8. ∎
Theorem 5.4.
For all primes $p$ such that $v_{p}(m)\neq v_{p}(n)$, we have
$p\mid 2\cdot\Delta_{F}\cdot\mathrm{N}(\mathfrak{f}_{E})\cdot\mathrm{N}(\mathfrak{f}_{E^{\prime}}).$
Proof.
The proof is the same as for Theorem 3.14, using Proposition 5.2 and Theorem 5.3 instead of Corollary 2.9 and Corollary 3.10. ∎
Theorem 5.5.
Suppose that $m$ is odd, $n\nmid m$, $\zeta_{n}\notin F$ and $\rho_{E,m}(G_{F})$ contains $\mathrm{SL}_{2}(m)$. Then
$3\mid m,\quad\text{and}\quad\mathrm{D}(\rho_{E,m}(G_{F}))\neq\mathrm{SL}_{2}(m),\quad\text{and}\quad F(\zeta_{n})\subseteq L$
with $L$ a $\mathbb{Z}/3\mathbb{Z}$-extension of $F(\zeta_{m})$.
Proof.
The proof is exactly the same as Theorem 4.4, replacing $F(E[n])$ by $F(E^{\prime}[n])$. ∎
Remark 5.6.
Except in Subsection 3.2, the results of Section 3 use the reduction link between $\rho_{E,p^{k+1}}$ and $\rho_{E,p^{k}}$ on the same elliptic curve. In particular, this method does not apply for different elliptic curves.
Appendix A Derived groups of $\mathrm{GL}_{2}(m)$ and $\mathrm{SL}_{2}(m)$
In this appendix, we give elementary and detailed proofs of well-known results about the derived groups of $\mathrm{GL}_{2}(m)$ and $\mathrm{SL}_{2}(m)$, for any integer $m$. They are used in Section 4 in the case where $m$ is odd. For a group $G$, we denote by $\operatorname{\mathrm{D}}(G)$ its commutator subgroup, generated by all the elements $[g,h]=ghg^{-1}h^{-1}$ with $g,h\in G$. We know that $\operatorname{\mathrm{D}}(G)$ is normal in $G$ and is the smallest group such that $G/\operatorname{\mathrm{D}}(G)$ is an abelian group: the abelianization of $G$.
We recall that $\mathrm{SL}_{2}(\mathbb{Z})$ is generated by $S=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$ and $T=\begin{pmatrix}1&1\\ 0&1\end{pmatrix}$. They satisfy $S^{2}=(ST)^{3}=-I$.
Proposition A.1.
The quotient group $\mathrm{SL}_{2}(\mathbb{Z})/\operatorname{\mathrm{D}}(\mathrm{SL}_{2}(\mathbb{Z}))$ is cyclic of order $12$, generated by the equivalence class of $T$.
Proof.
Let $\bar{S}$ and $\bar{T}$ the classes of $S$ and $T$ in $\mathrm{SL}_{2}(\mathbb{Z})/\mathrm{D}(\mathrm{SL}_{2}(\mathbb{Z}))$ respectively. Since $\mathrm{SL}_{2}(\mathbb{Z})/\mathrm{D}(\mathrm{SL}_{2}(\mathbb{Z}))$ is abelian, we have $\bar{S}\bar{T}=\bar{T}\bar{S}$. Hence $(\bar{S}\bar{T})^{3}=\bar{S}^{3}\bar{T}^{3}=\bar{S}^{2}$, which gives $\bar{S}=\bar{T}^{-3}$ and $\bar{T}^{12}=\bar{S}^{4}=I$. It follows that $\bar{T}$ generated $\mathrm{SL}_{2}(\mathbb{Z})/\mathrm{D}(\mathrm{SL}_{2}(\mathbb{Z}))$ and has order $12$. ∎
Let $m\mid n$ be positive integers and, for $i=n,m$, set $X_{i}=\mathrm{SL}_{2}(i))/\mathrm{D}(\mathrm{SL}_{2}(i))$. The following diagram has exact rows and is commutative.
[…]
Proposition A.2.
Let $m$ be a positive integer. The abelianization of $\mathrm{SL}_{2}(m)$ is isomorphic to $\mathbb{Z}/\gcd(m,12)\mathbb{Z}$. In particular, if $m$ is coprime to $6$, then $\mathrm{SL}_{2}(m)$ is perfect.
Proof.
By Diagram (5) and Proposition A.1, the image of $T$ in $\mathrm{SL}_{2}(m)$ generates the abelianization of $\mathrm{SL}_{2}(m)$, and its order divides $12$. Moreover, the image of $T$ in $\mathrm{SL}_{2}(m)$ has order $m$. Hence the abelianization of $\mathrm{SL}_{2}(m)$ has order dividing $\gcd(m,12)$. To prove that its order is exactly $\gcd(m,12)$, it suffices to prove it for $m=2,3$ and $4$, and then to use again Diagram (5). We define the commutators
$A=\left[\begin{pmatrix}-1&-1\\ 0&-1\end{pmatrix},\begin{pmatrix}1&-2\\ 1&-1\end{pmatrix}\right]=\begin{pmatrix}3&-1\\ 1&0\end{pmatrix},$
$B=\left[\begin{pmatrix}-1&-1\\ 0&-1\end{pmatrix},\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\right]=\begin{pmatrix}2&1\\ 1&1\end{pmatrix}.$
Let $G=\langle A,B\rangle\subseteq\mathrm{SL}_{2}(\mathbb{Z})$. For $m=2,3,4$, let $G_{m}$ be the image of $G$ is $\mathrm{SL}_{2}(3)$. By computing explicitely $G_{2}$, $G_{3}$ and $G_{4}$, we find that they have respectively order $3$, $8$ and $12$, so $\mathrm{SL}_{2}(m)/G_{m}$ has order $m$ and is abelian. Hence, as a subgroup of $\mathbb{Z}/12\mathbb{Z}$, it is isomorphic to $\mathbb{Z}/m\mathbb{Z}=\mathbb{Z}/\gcd(m,12)\mathbb{Z}$. ∎
Remark A.3.
We also can prove the previous result for $m=p^{k}$ with $p\geq 5$ prime, just by observing that the diagram  5 gives that the order of $X_{m}$ is both a divisor of $12$ and a divisor of #$\#\mathrm{SL}_{2}(p^{k})=p^{3(k-1)+1}(p-1)(p+1)$. Hence, the order of the abelianization of $\mathrm{SL}_{2}(p^{k})$ is $1$, and so its derived group is itself, unless $p=2$ or $3$.
Proposition A.4.
For an odd integer $m$, the derived group of $\mathrm{GL}_{2}(m)$ is $\mathrm{SL}_{2}(m)$, and so its abelianization is isomorphic to $(\mathbb{Z}/m\mathbb{Z})^{*}$. If $m$ is even, the derived group of $\mathrm{GL}_{2}(m)$ has index two in $\mathrm{SL}_{2}(m)$.
Proof.
For $A,B\in\mathrm{GL}_{2}(m)$, we have $\det(ABA^{-1}B^{-1})=1$. So $[A,B]\in\mathrm{SL}_{2}(m)$. Therefore
$\operatorname{\mathrm{D}}(\mathrm{SL}_{2}(m))\leq\operatorname{\mathrm{D}}(\mathrm{GL}_{2}(m))\leq\mathrm{SL}_{2}(m).$
For $m$ coprime to $6$, we have proven that $\operatorname{\mathrm{D}}(\mathrm{SL}_{2}(m))=\mathrm{SL}_{2}(m)$, and so $\operatorname{\mathrm{D}}(\mathrm{GL}_{2}(m))=\mathrm{SL}_{2}(m)$. For $m=3^{k}$, we know that $\mathrm{D}(\mathrm{SL}_{2}(3^{k}))$ has index $3$ in $\mathrm{SL}_{2}(3^{k})$, and so $\mathrm{D}(\mathrm{GL}_{2}(3^{k}))$ is either $\mathrm{SL}_{2}(3^{k})$ or $\mathrm{D}(\mathrm{SL}_{2}(3^{k}))$. For $k=1$, $\mathrm{D}(\mathrm{SL}_{2}(3))$ is explicitly know by the proof of Proposition A.2 and
$\left[\begin{pmatrix}1&-1\\ 1&1\end{pmatrix},\begin{pmatrix}1&1\\ 0&-1\end{pmatrix}\right]=\begin{pmatrix}1&1\\ -1&0\end{pmatrix}\notin\mathrm{D}(\mathrm{SL}_{2}(3)).$
It follows that $\mathrm{D}(\mathrm{GL}_{2}(3))=\mathrm{SL}_{2}(3)$. We obtain the diagram below:
[…]
showing that $\operatorname{\mathrm{D}}(\mathrm{SL}_{2}(3^{k}))\neq\operatorname{\mathrm{D}}(\mathrm{GL}_{2}(3^{k}))$ and so $\operatorname{\mathrm{D}}(\mathrm{GL}_{2}(3^{k}))=\mathrm{SL}_{2}(3^{k})$. For $m=2^{k}$, we use the same strategy. From Proposition A.2, we already know that $\operatorname{\mathrm{D}}(\mathrm{GL}_{2}(2))$ has index $2$ in $\mathrm{SL}_{2}(2)$. We have the following diagram:
[…]
The group $\operatorname{\mathrm{D}}(\mathrm{SL}_{2}(4))$ is known by the proof of Proposition A.2, and we have
$\left[\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\begin{pmatrix}2&1\\ -1&1\end{pmatrix}\right]=\begin{pmatrix}0&-1\\ 1&1\end{pmatrix}\in\operatorname{\mathrm{D}}(\mathrm{GL}_{2}(4))\backslash\operatorname{\mathrm{D}}(\mathrm{SL}_{2}(4)).$
Using that $\operatorname{\mathrm{D}}(\mathrm{SL}_{2}(2^{k}))$ has index $4$ in $\mathrm{SL}_{2}(2^{k})$ for $k\geq 2$, we obtain that $\operatorname{\mathrm{D}}(\mathrm{GL}_{2}(4))$ has exactly index $2$ in $\mathrm{SL}_{2}(4)$ and $\operatorname{\mathrm{D}}(\mathrm{GL}_{2}(2^{k}))$ has exactly index $2$ in $\mathrm{SL}_{2}(2^{k})$ for all $k\geq 2$. The result follows, since the Chinese remainder theorem gives
$\operatorname{\mathrm{D}}(\mathrm{GL}_{2}(m))\simeq\underset{\underset{p\text{prime}}{p^{k}\mid\mid m}}{\prod}\operatorname{\mathrm{D}}(\mathrm{GL}_{2}(p^{k})).$